### v2

# Azumaya algebras and Barr’s Theorem

Azumaya algebras and Barr’s Theorem
Thierry Coquand Computer Science and Engineering Department, University of Gothenburg, Sweden, coquand@chalmers.se Henri Lombardi Université Marie et Louis Pasteur, CNRS, LmB, FR-25000 Besançon, France, henri.lombardi@univ-fcomte.fr, stefan.neuwirth@univ-fcomte.fr Stefan Neuwirth Université Marie et Louis Pasteur, CNRS, LmB, FR-25000 Besançon, France, henri.lombardi@univ-fcomte.fr, stefan.neuwirth@univ-fcomte.fr
(May 17, 2025)
Introduction
The goal of this paper is to show how one can use the notion of classifying topos of a coherent theory [Joh, CLR, Wraith], here the theory of strictly Henselian local rings [Wraith], to prove some basic results about Azumaya algebras in a constructive setting.
A fundamental invariant associated to any commutative field $F$ is its Brauer group, whose elements are division algebras over $F$, up to isomorphism. Alternatively an element of the Brauer group can be described as a central simple algebra, up to Morita equivalence, and the group operation is then the tensor product. Two important characterisations of the notion of central simple algebra are the following:
  1. 1.
    an algebra $A$ over $F$ which is finite and such that the canonical map $A\otimes_{F}A^{op}\rightarrow End_{F}(A)$ is an isomorphism; or, equivalently,
  2. 2.
    an algebra $A$ which becomes a matrix algebra over an algebraic (resp. separable) extension of $F$.
This notion of central simple algebra over a field has been generalized to the notion of Azumaya algebra over an arbitrary commutative ring $R$. The definition is a generalisation of the first characterisation: an algebra $A$ which is projective of finite type over $R$ and such that the canonical map $A\otimes_{F}A^{op}\rightarrow End_{R}(A)$ is an isomorphism. In this note, we show in a constructive setting that this definition is equivalent to a suitable generalisation of the second characterisation, using a constructively valid version of Barr’s Theorem [Joh].
1 Some coherent theories
We will consider some coherent theories [Joh] in the language of rings.
1.1 Theory of local rings
The theory of local rings can be formulated in a coherent way with the axioms
$0\neq 1~{}~{}~{}~{}~{}~{}~{}~{}~{}\mathsf{inv}(x)\vee\mathsf{inv}(1-x)$
where $\mathsf{inv}(x)$ denotes $\exists_{y}~{}1=xy$.
If we add to our language a new predicate $\mathsf{m}(x)$ with the axioms
$\mathsf{inv}(x)\vee\mathsf{m}(x)~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\neg(\mathsf{m}(x)\wedge\mathsf{inv}(x))$
then we get the theory of local rings that are residually discrete, i.e. where the property of being invertible is now decidable, and the predicate $\mathsf{m}(x)$ corresponds to the unique maximal ideal of the ring. In this theory, we can express in a coherent way that the ring is Henselian by stating that for any monic polynomial $P$, if we have $u$ such that $P(u)\in\mathsf{m}$ and $P^{\prime}(u)\notin\mathsf{m}$, we can find $\alpha$ such that $P(\alpha)=0$ and $\alpha-u\in\mathsf{m}$. Equivalently [ALP], any monic polynomial of degree $l+1$ which is of the form $X^{l}(X-1)$ mod. $\mathsf{m}$ has a root which is equal to $1$ mod. $\mathsf{m}$. The work [ALP] shows in an intuitionistic setting for this theory the lifting property for idempotents of a finitely generated algebra over a Henselian local ring.
It does not seem possible however to express the property of being Henselian in a coherent way in the language of rings (i.e. without introducing the predicate $\mathsf{m}(x)$). Surprisingly, as noticed by G. Wraith [Wraith], the property of being strictly Henselian, corresponding classically to the property of being Henselian with the residue field being separably closed, can be expressed in a coherent way. In this note, we present some remarks connected to this observation. The first one is a small variation of Wraith’s axiomatisation. Both axiomatisations can be used to define the étale site over a ring. Finally, we show in a constructive setting that Azumaya algebras over a ring are exactly the algebras that are locally matrix algebras for the étale topology.
1.2 Theory of separably closed local rings
If $R$ is a ring, we define the Zariski spectrum of $R$ to be the distributive lattice generated by symbols $D(a)$ and relations
$D(1)=1~{}~{}~{}~{}~{}D(0)=0~{}~{}~{}~{}~{}~{}D(a+b)\leqslant D(a)\vee D(b)~{}~{}~{}~{}~{}~{}~{}~{}D(ab)=D(a)\wedge D(b)$
We write $D(a_{1},\dots,a_{n})$ for $D(a_{1})\vee\dots\vee D(a_{n})$.
One way to build this lattice is to define $D(a_{1},\dots,a_{n})$ to be the radical of the ideal $(a_{1},\dots,a_{n})$.
In particular, this shows that $D(a_{1},\dots,a_{n})=1$ if, and only if, $(a_{1},\dots,a_{n})=1$.
It is also direct to see that $D(a,b)=D(a+b,ab)$ and more generally that
$D(a_{1},\dots,a_{n})=D(s_{1},\dots,s_{n})\text{,}$
where $s_{1},\dots,s_{n}$ are the elementary symmetric functions of $a_{1},\dots,a_{n}$.
For the formulation of this theory of strictly Henselian rings, we need the notion of universal decomposition algebra $L=R[x_{1},\dots,x_{n}]=R[X_{1},\dots,X_{n}]/I$ of a monic polynomial
$f=X^{n}-u_{1}X^{n-1}+u_{2}X^{n-2}-\dots$
in $R[X]$, where $I$ is the ideal generated by $u_{1}-\sigma_{1},~{}\dots,~{}u_{n}-\sigma_{n}$, where $\sigma_{1},\dots,\sigma_{n}$ in $R[X_{1},\dots,X_{n}]$ are the elementary symmetric polynomials. We know that $L=R[x_{1},\dots,x_{n}]$ is freely generated as a module over $R$ and it follows from this that $L$ is a faithfully flat extension of $R$.
We define $\delta_{i}(f)$ to be the element in $R$ equal to $\sigma_{i}(f^{\prime}(x_{1}),\dots,f^{\prime}(x_{n}))$ and $\Delta(f)$ to be $D(\delta_{1}(f),\dots,\delta_{n}(f))$.
We say that $f$ is unramifiable if, and only if, $1=\Delta(f)$. Since $D(f^{\prime}(x_{1}),\dots,f^{\prime}(x_{n}))=D(\delta_{1}(f),\dots,\delta_{n}(f))$, this condition is equivalent to $(f^{\prime}(x_{1}),\dots,f^{\prime}(x_{n}))=1$ in $L$.
Over a residually discrete local ring with residue field $F$, to be unramifiable means that the polynomial $f$ has a simple root in some extension of $F$.
Lemma 1.1.
Let $R$ be a local ring residually discrete of maximal ideal $\mathsf{m}$ with residue field $F=R/\mathsf{m}$. If $f$ is a monic polynomial in $R[X]$ which has a simple root in some nontrivial $F$-algebra $A$, then $f$ is unramifiable.
Lemma 1.2.
If $f=(X-a_{1})\dots(X-a_{p})g$ in $R[X]$ with $g$ monic of degree $n$ and $\Delta(f)=1$ and $R$ is local, then $f^{\prime}(a_{1})$ is invertible or $\dots$ or $f^{\prime}(a_{p})$ is invertible or $\Delta(g)=1$.
Proof.
We write $h=(X-a_{1})\cdots(X-a_{p})$.
We have in $R[x_{1},\dots,x_{n}]$, the decomposition algebra of $g$,
$1=D(f^{\prime}(a_{1}),\dots,f^{\prime}(a_{p}),h(x_{1})g^{\prime}(x_{1}),\dots,h(x_{n})g^{\prime}(x_{n}))$
and hence
$1=D(f^{\prime}(a_{1}),\dots,f^{\prime}(a_{p}),g^{\prime}(x_{1}),\dots,g^{\prime}(x_{n}))$
and from this follows
$1=D(f^{\prime}(a_{1}),\dots,f^{\prime}(a_{p}))\vee\Delta(g)$
in $R[x_{1},\dots,x_{n}]$. Since the decomposition algebra is faithfully flat, we get
$1=D(f^{\prime}(a_{1}),\dots,f^{\prime}(a_{p}))\vee\Delta(g)$
in $R$, and since $R$ is local we have the conclusion. ∎
Lemma 1.3.
If $R$ is local and any monic unramifiable polynomial has a root in $R$ then any monic unramifiable polynomial has a simple root in $R$.
Proof.
Assume that $R$ is local and any monic unramifiable polynomial has a root in $R$. Let $f$ be a monic unramifiable polynomial. It has a root $a_{1}$ and we can write $f=(X-a_{1})f_{1}$. By Lemma 1.2, $a_{1}$ is a simple root of $f$ or $\Delta(f_{1})=1$. If $\Delta(f_{1})=1$ then it has a root $a_{2}$ and $f=(X-a_{1})(X-a_{2})f_{2}$. Then by Lemma 1.2 again, $a_{1}$ or $a_{2}$ is a simple root of $f$ or $\Delta(f_{2})=1$, and so on until we find a simple root of $f$. ∎
Let $k$ be a ring. We consider now the category of finitely presented $k$-algebras with the covering
  1. 1.
    $S$ is covered by $S\rightarrow S[1/u_{1}],\dots,S\rightarrow S[1/u_{m}]$ if $1=(u_{1},\dots,u_{m})$;
  2. 2.
    $S$ is covered by $S\rightarrow S[X]/(P)$ if $P$ is monic and unramifiable.
It also follows from Lemma 1.3 that we can instead take the following covering, since it classifies an equivalent theory.
  1. 1.
    $S$ is covered by $S\rightarrow S[1/u_{1}],\dots,S\rightarrow S[1/u_{m}]$ if $1=(u_{1},\dots,u_{m})$;
  2. 2.
    $S$ is covered by $S\rightarrow S[X]/(P)[1/P^{\prime}]$ if $P$ is monic and unramifiable.
This defines a site, in the sense of Grothendieck (see e.g. [Joh]), over the opposite of the category of finitely presented $k$-algebras. The corresponding sheaf model is the classifying topos of the theory of strictly Henselian $k$-algebras [Joh]. In this sheaf model, we have a “generic” $k$-algebra which is strictly Henselian.
We also may state, following Wraith,
Theorem 1.1.
If $R$ is a local ring which is also residually discrete, then $R$ is strictly Henselian if, and only if, $R$ is Henselian and its residue field is separably closed.
Proof.
Let $P$ be a monic polynomial which has residually a simple root $u$. We show by induction on the degree of $P$ that $P$ has a root in $R$ which is equal to $u$ mod. $\mathsf{m}$. The polynomial $P$ is unramifiable by Lemma 1.1, since it has residually a simple root, and it has a root $a$. If $a\neq u$ mod. $\mathsf{m}$ then we write $P=(X-a)Q$, and $u$ is residually a simple root of $Q$. ∎
2 Some lemmas about Henselian rings
In the next section, we need some results about Henselian rings. In order to make the paper self-contained we list here these results with their proofs. They are taken from the reference [ALP], with a variation in the proof of Lemma 2.1.
We assume in this section that $R$ is local residually discrete with maximal ideal $\mathsf{m}$ and Henselian.
Lemma 2.1 ([ALP, Prop. 5.1]).
Let $P$ be a monic polynomial of degree $n$ in $R[X]$ and $R[x]=R[X]/(P)$. If we have $e(X)$ in $R[X]$ such that $e(x)$ is idempotent mod. $\mathsf{m}R[x]$, then we can find $u$ idempotent in $R[x]$ and $u=e(x)$ mod. $\mathsf{m}R[x]$.
Proof.
Let $A=R[x_{1},\dots,x_{n}]$ be the universal decomposition algebra of $P$. This is a free $R$-module over $R$ of dimension $n!$ and the symmetric group acts on it [LQ, Fact 3-4.4]. We may identify $R[x]$ with the subalgebra $R[x_{1}]$ and we have $e(x_{1})$ in $R[x_{1}]$ which is idempotent mod. $\mathsf{m}R[x_{1}]$. We show that we have $u_{1}$ in $R[x_{1}]$ which is idempotent in $R[x_{1}]$ and equal to $e(x_{1})$ mod. $\mathsf{m}R[x_{1}]$.
The product $e(x_{1})\cdots e(x_{n})$ is in $R$ and it is an idempotent mod. $\mathsf{m}$, so it is $1$ or 00 mod. $\mathsf{m}$. If it is $1$, we have $e(x_{1})=1$ mod. $\mathsf{m}R[x_{1}]$ and we can take $u_{1}=1$.
We can thus assume $e(x_{1})\cdots e(x_{n})=0$ mod. $\mathsf{m}$. We write $e_{I}=\prod_{i\in I}e(x_{i})$ for $I$ subset of $1,\dots,n$. We let [^1] The $e_{I},~{}|I|=k$, are pairwise conjugate by the action of the symmetric group and equality is decidable mod. $\mathsf{m}$. $k<n$ be such that $e_{I}\neq 0$ mod. $\mathsf{m}$ if $|I|=k$ and $e_{I}=0$ mod. $\mathsf{m}$ if $|I|=k+1$. Note that the $e_{I}$ for $|I|=k$ are pairwise orthogonal mod. $\mathsf{m}$. We have $\Sigma_{|I|=k}e_{I}=1$ mod. $\mathsf{m}$ since it is an element in $R$, which is idempotent mod. $\mathsf{m}$. Hence the family of the $e_{I},~{}|I|=k$, forms an FSOI [^2] A “Fundamental System of Orthogonal Idempotents” ([LQ, Sec. II-4]). mod. $\mathsf{m}A$.
We deduce that
$Q(X)=\prod_{|I|=k}(X-e_{I})\text{,}$
which is a monic polynomial in $R[X]$, is of the form $X^{N-1}(X-1)$ mod. $\mathsf{m}R[X]$ with $N=\binom{n}{k}$.
Since $R$ is Henselian, we have $\alpha$ in $R$ such that $Q(\alpha)=0$ and $\lambda=Q^{\prime}(\alpha)$ is invertible. If we write $Q_{I}=Q/(X-e_{I})$ then $Q^{\prime}(X)=\Sigma_{I}Q_{I}$. We take
$v_{I}=\frac{1}{\lambda}Q_{I}(\alpha)$
and we then have that the $v_{I},~{}|I|=k$, form an FSOI in $R[x_{1},\dots,x_{n}]$. We have $v_{I}=\Pi_{J\neq I}(1-e_{J})=e_{I}$ mod. $\mathsf{m}A$. Let $L_{1}$ be the set of all $I$ such that $1$ is in $I$ and $|I|=k$. If we write $u_{1}$ for $\Sigma_{I\in L_{1}}v_{I}$, we have $u_{1}$ in $R[x_{1}]$ and $u_{1}$ is equal to $\Sigma_{I\in L_{1}}e_{I}=e(x_{1})$ mod. $\mathsf{m}A$. This equality follows from $e(x_{1})=e(x_{1})(\Sigma_{I}e_{I})=\Sigma_{I}e(x_{1})e_{I}=\Sigma_{I\in L_{1}}e_{I}$ mod. $\mathsf{m}A$. Hence $e(x_{1})-u_{1}\in R[x_{1}]\cap\mathsf{m}A=\mathsf{m}R[x_{1}]$. ∎
For completeness, we recall the statement and proof of Nakayama’s Lemma [LQ, Lemma IX-2.1].
Lemma 2.2.
Let $M$ be a finitely generated $R$-module and $N$ an $R$-submodule such that $M=N+\mathsf{m}M$, then $M=N$.
Proof.
Let $u$ be a vector column of generators of $M$, then we can find a vector $v$ in $N$ and a matrix $P$ with elements in $\mathsf{m}$ such that $u=v+Pu$. We then have $(I-P)u=v$ and hence the elements of $v$ are generators of $M$ since $\det(I-P)$ is invertible. ∎
Corollary 2.1.
Let $A$ be an $R$-algebra which is finitely generated as an $R$-module, and $u$ in $A$ which is invertible mod. $\mathsf{m}A$, then $u$ is invertible in $A$.
Proof.
We have $A=uA+\mathsf{m}A$ and we can apply the previous Lemma. ∎
Lemma 2.3 ([ALP, Prop. 5.8]).
Let $P$ in $R[X]$ be a monic polynomial in $R[X]$ and assume that we have a decomposition $P=fg$ mod. $\mathsf{m}R[X]$ with $f$ and $g$ monic polynomials coprime mod. $\mathsf{m}$. Then we can find a decomposition $P=FG$ in $R[X]$ with $F=f$ and $G=g$ mod. $\mathsf{m}$.
Proof.
We let $R[x]=R[X]/(P)$. We can find an idempotent $e$ in $R[x]$ mod. $\mathsf{m}R[x]$ such that $(f)=(e),~{}(g)=(1-e)$ mod. $\mathsf{m}R[x]$ [ALP, Prop. 3.11]. Using Lemma 2.1, we may assume $e$ idempotent in $R[x]$. Let $n$ (resp. $m$) be the degree of $f$ (resp. $g$). Then $e,ex,\dots,ex^{m-1}$ and $(1-e),(1-e)x,\dots,(1-e)x^{n-1}$ generate $R[x]$ as an $R$-module mod. $\mathsf{m}$. By Lemma 2.2, they generate the $R$-module $R[x]$. In particular, $ex^{m}$ belongs to the $R$-module generated by $e,ex,\dots,ex^{m-1}$ (since $e$ is idempotent in $R[x]$) and we find in this way $G$ monic of degree $m$ such that $eG(x)=0$ in $R[x]$. Similarly we find $F$ monic of degree $n$ such that $(1-e)F(x)=0$ in $R[x]$. We then have $F(x)G(x)=F(x)eG(x)=0$ and hence $P$ divides $FG$. But they have the same degree, hence $P=FG$. ∎
Theorem 2.1 ([ALP, Thm. 5.9]).
If $A$ is an $R$-algebra which is finitely generated as an $R$-module and $a$ in $A$ is idempotent mod. $\mathsf{m}A$, then we can find $e$ in $A$ idempotent in $A$ and such that $e=a$ mod. $\mathsf{m}A$.
Proof.
We consider the subalgebra $R[a]$ of $A$. Since $A$ is finitely generated as an $R$-module we first find $F$ monic in $R[X]$ such that $F(a)=0$. We decompose $F=X^{n}(1-X)^{m}h$ mod. $\mathsf{m}R[X]$, with $h$ monic and $h(0),~{}h(1)$ not in $\mathsf{m}$. If $n=0$ we have $F(0)=h(0)$ mod. $\mathsf{m}$. Since $F(a)=0$, we get that $a$ is invertible mod. $\mathsf{m}A$, and hence invertible in $A$ using Corollary 2.1, and we can take $e=1$. Similarly, if $m=0$, we get that $1-a$ is invertible and we can take $e=0$.
In the remaining case, we can find using Lemma 2.3 a decomposition in monic polynomials $F=PQH$ in $R[X]$ with $P=X^{n},~{}Q=(1-X)^{m},~{}H=h$ mod. $\mathsf{m}$. We also have $U$ and $V$ in $R[X]$ such that $X(1-X)U+HV=Res(X(1-X),H)=1$ [LQ, Fact 3-7.1], [ALP, Lemma 3.10], and so $H(a)V(a)=1$ mod. $\mathsf{m}A$, and so $H(a)$ is invertible mod. $\mathsf{m}A$ and so invertible in $A$ using Corollary 2.1. Since $F(a)=P(a)Q(a)H(a)=0$, we have $P(a)Q(a)=0$ in $A$. We also have $P(a)=a$ and $Q(a)=1-a$ mod. $\mathsf{m}A$ and so $P(a)+Q(a)=1$ mod. $\mathsf{m}A$ and is also invertible in $A$, by Corollary 2.1, with inverse $\mu$. We can then take $e=\mu P(a)$. ∎
3 Azumaya algebras
An Azumaya algebra $A$ over a ring $k$ is an algebra such that
  1. 1.
    $A$ is projective of finite type;
  2. 2.
    the canonical map $A\otimes_{R}A^{op}\rightarrow End_{k}(A)$ is an isomorphism.
Note that we allow the trivial algebra to be an Azumaya algebra.
We have the following characterisation, which generalizes the result that an algebra is central simple over a field if, and only if, it can be split by a separable extension of this field.
Theorem 3.1.
$A$ is an Azumaya algebra over $k$ if, and only if, we can build a finite tree of root $k$ using the following, where all leaves are matrix algebras:
  1. 1.
    $S$ is covered by $S\rightarrow S[1/u_{i}]$ if $1=(u_{1},\dots,u_{n})$;
  2. 2.
    $S$ is covered by $S\rightarrow S[X]/(P)$ if $P$ is monic and unramifiable.
If we have such a tree starting from $k$, notice that each extension $S\rightarrow S[X]/(P)$ for $P$ monic is faithfully flat, and $S\rightarrow\prod_{i}S[1/u_{i}]$ if $1=(u_{1},\dots,u_{n})$ is also faithfully flat. It follows then that $A$ is finite projective [LQ, Thm. VIII-6.7]. The map $A\otimes_{k}A^{op}\rightarrow End_{k}(A)$ is then a bijection since it becomes a bijection by faithfully flat extension.
The converse follows from the following two results.
Lemma 3.1.
If $A$ is an Azumaya algebra over $R$, which is a local and residually discrete and separably closed $k$-algebra, then $A$ is a matrix algebra.
Proof.
We know that $R$ is Henselian and that $R/\mathsf{m}$ is separably closed by Theorem 1.1.
We simplify the argument in Milne’s course notes ([Milne, Chapter IV, Proposition 1.6]), using the constructive development in [ALP] and [CLN]. If $\mathsf{m}$ is the maximal ideal of $R$, we have that $A$ is split over $R/\mathsf{m}$ ([CLN]), and hence we have a matrix algebra decomposition $e_{ij}$ mod. $\mathsf{m}A$. We can then lift the idempotent $e_{11}$ mod. $\mathsf{m}A$ to some idempotent $e$ in $R$ using Theorem 2.1. We have that $Ae$ is projective and hence free since $R$ is a local ring. The map $A\rightarrow End_{R}(Ae)$ is then an isomorphism, since it is a map between two free $R$-modules of the same dimension, and it is residually an isomorphism. ∎
Corollary 3.1.
If $A$ is an Azumaya algebra over $R$, which is a local and separably closed $k$-algebra, then $A$ is a matrix algebra.
Proof.
For a given Azumaya algebra, the conclusion is coherent, and the hypothesis that $A$ is an Azumaya algebra is formulated in a coherent theory. Since it is proved with the extra hypothesis that $R$ is residually discrete, the syntactic version of Barr’s Theorem [^3] We can actually follow the argument in [Joh] and apply Friedman-Dragalin’s translation [Avi] w.r.t. to the proposition expressing that $A$ is a matrix algebra. (constructively valid) shows that it can be proved without this hypothesis. ∎
Theorem 3.1 follows by interpreting this result in the classifying topos of the theory of strictly Henselian $k$-algebra.
The statement of Theorem 3.1 can be seen as a generalisation of the fact that over a field, a central simple algebra becomes a matrix algebra by a faithfully flat extension [CLN] (classically over an algebraic extension).
We can use such a tree to give arguments by “tree induction”. We give some simple examples.
Lemma 3.2.
Any Azumaya algebra $A$ is central.
Proof.
This means that if $a$ in $A$ satisfies $ax=xa$ for all $x$ in $A$ then $a$ is in $R$. We prove this by “tree induction” using Theorem 3.1. The statement holds on the leaves (it holds for a matrix algebra) and descends to the root. ∎
Lemma 3.3.
If $R$ is a local ring and $A$ a matrix algebra of rank $>0$ [LQ, Def. X-2.2] then any automorphism of $A$ is of the form $x\mapsto axa^{-1}$ for some $a$ in $GL_{n}(A)$.
Proof.
Let $A=End_{R}(R^{n})$ with $v_{1},\dots,v_{n}$ the canonical basis of $R^{n}$ and define $e_{ij}v_{k}=\delta_{jk}v_{i}$. Let $\psi$ be an automorphism of $A$. If $p_{ij}=\psi(e_{ij})$, we have that the $p_{i}=p_{ii}$ form an FSOI and $p_{ij}p_{kl}=\delta_{jk}p_{il}$. Since $R$ is local, each $V_{i}=p_{i}(R^{n})$ is free. Furthermore $R^{n}$ is a direct sum of the $V_{i}$’s; the $V_{i}$’s are pairwise isomorphic since $p_{ji}$ defines an isomorphism between $V_{i}$ and $V_{j}$. So they are all free of rank $1$. Let $w_{1}$ be a basis of $V_{1}$ and $w_{j}=p_{j1}w_{1}$. If $a(v_{i})=w_{i}$ we have $\psi(x)=axa^{-1}$. ∎
Theorem 3.2 (Skolem-Noether).
If $A$ is an Azumaya algebra, of rank $>0$ as a projective module over a ring $k$, and $\psi$ is an automorphism of $A$, then $M=\{a\in A~{}|~{}\forall_{x}ax=\psi(x)a\}$ is a projective $k$-module of rank $1$.
Proof.
We use Theorem 3.1. The statement holds on the leaves ($M$ is even free of rank $1$ on the leaves by Lemma 3.2) and descends to the root by tree induction using the local-global principle and the descent principle w.r.t. faithfully flat extension for projective modules ([LQ, Thm. VIII-6.7]). ∎
If $k$ is local, the module $M$ is free of rank $1$ and any generator $u$ of $M$ is invertible (since it is invertible on the leaves) and $\psi$ is the inner automorphism defined by $u$, so we recover the usual form of the Skolem-Noether Theorem (any automorphism of a central simple algebra is an inner automorphism).
Claude Quitté suggested the following refinement.
Theorem 3.3.
If $A$ is an Azumaya algebra, of rank $>0$ as a projective module over a ring $k$, and $\psi$ is an automorphism of $A$, then
$M=\{a\in A~{}|~{}\forall_{x}ax=\psi(x)a\}$ and $N=\{a\in A~{}|~{}\forall_{x}a\psi(x)=xa\}$
are projective $k$-modules of rank $1$ and $M\otimes_{k}N\simeq k$ and $MN=k1$. Furthermore, $M$ and $N$ are direct factors in $A$.
Proof.
We have on the leaves $MN\subseteq k1$ and this statement descends to the root by tree induction. This can be used to define a map $M\rightarrow Hom(N,k),~{}x\mapsto(y\mapsto xy)$. This map is an isomorphism on the leaves, and this also descends to the root by tree induction. Finally, $M$ and $N$ are direct factors on the leaves, and this fact descends as well. ∎
4 Theory of algebraically closed local rings
To get the theory of algebraically closed local rings, we add to the theory of local rings the axioms
$\exists_{x}~{}P(x)=0$
for any monic nonconstant polynomial $P$.
Wraith conjectured in [Wraith] that this should define the classifying theory of the fppf topos. This question is discussed further in Blechschmidt’s PhD thesis [BT].
The same argument as in Lemma 3.1 shows the following result.
Lemma 4.1.
If $A$ is an Azumaya algebra over $R$, which is a local and residually discrete and algebraically closed $k$-algebra, then $A$ is a matrix algebra.
We can then prove as in the previous section
Theorem 4.1.
$A$ is an Azumaya algebra over $k$ if, and only if, we can build a finite tree of root $k$ using the following, where all leaves are matrix algebras:
  1. 1.
    $S$ is covered by $S\rightarrow S[1/u_{i}]$ if $1=(u_{1},\dots,u_{n})$;
  2. 2.
    $S$ is covered by $S\rightarrow S[X]/(P)$ if $P$ is monic nonconstant.
5 Conclusion
This paper is a beginning of the theory of Azumaya algebras in a constructive setting. The natural next step will be to analyse Gabber’s Theorem [KO] that the group of Azumaya algebras is the torsion part of the second cohomology group with coefficients in $GL_{n}(R)$ for the étale topology.

### v3

@@
 Azumaya algebras and Barr’s Theorem
-Thierry Coquand Computer Science and Engineering Department, University of Gothenburg, Sweden, coquand@chalmers.se Henri Lombardi Université Marie et Louis Pasteur, CNRS, LmB, FR-25000 Besançon, France, henri.lombardi@univ-fcomte.fr, stefan.neuwirth@univ-fcomte.fr Stefan Neuwirth Université Marie et Louis Pasteur, CNRS, LmB, FR-25000 Besançon, France, henri.lombardi@univ-fcomte.fr, stefan.neuwirth@univ-fcomte.fr
-(May 17, 2025)
+Thierry Coquand Henri Lombardi Stefan Neuwirth
+(November 24, 2025)
 Introduction
@@
 Alternatively an element of the Brauer group can be described as a central simple algebra, up to Morita equivalence, and the group operation is then the tensor product.
-Two important characterisations of the notion of central simple algebra are the following:
+Two important characterisations of the notion of central simple algebra are the following [^1] We study these characterisations in a constructive setting in the reference [CLN].:
   1. 1.
@@
   2. 2.
-    an algebra $A$ which becomes a matrix algebra over an algebraic (resp. separable) extension of $F$.
+    an algebra $A$ which becomes a matrix algebra over an algebraic extension of $F$, or, equivalently, which becomes a matrix algebra over an algebraic and separable extension of $F$.
 This notion of central simple algebra over a field has been generalized to the notion of Azumaya algebra over an arbitrary commutative ring $R$.
@@
 $\mathsf{inv}(x)\vee\mathsf{m}(x)~~~~~~~~~~\neg(\mathsf{m}(x)\wedge\mathsf{inv}(x))$
-then we get the theory of local rings that are residually discrete, i.e.
+then we get the theory of local rings that are residually discrete [^2] The terminology “discrete” for a set with a decidable equality, while conflicting with some topological intuitions, is well established in constructive mathematics since the work of Bishop [Bish], and we follow this usage in this paper., i.e.
 where the property of being invertible is now decidable, and the predicate $\mathsf{m}(x)$ corresponds to the unique maximal ideal of the ring.
@@
 1.2 Theory of separably closed local rings
-If $R$ is a ring, we define the Zariski spectrum of $R$ to be the distributive lattice generated by symbols $D(a)$ and relations
+If $R$ is a ring, and $a_{1},\dots,a_{n}$ elements of $R$, we write $(a_{1},\dots,a_{n})$ for the ideal of $R$ generated by $a_{1},\dots,a_{n}$.
+We define, following Joyal [Joy], the Zariski spectrum of $R$ as the universal support $D:R\rightarrow Z(R)$ over $R$.
+A support is a map $d:R\rightarrow M$ into a bounded distributive lattice $M$ satisfying the relations
 $d(1)=1~~~~~d(0)=0~~~~~~d(a+b)\leqslant d(a)\vee d(b)~~~~~~~~d(ab)=d(a)\wedge d(b)$
 We write $d(a_{1},\dots,a_{n})$ for $d(a_{1})\vee\dots\vee d(a_{n})$.
-One way to build this lattice is to define $D(a_{1},\dots,a_{n})$ to be the radical of the ideal $(a_{1},\dots,a_{n})$.
+It is direct that $d(a)\leqslant d(a_{1},\dots,a_{n})$ as soon as $a$ belongs to the radical of the ideal $(a_{1},\dots,a_{n})$.
+Conversely, it is elementary to check that we get a support if we define $D(a_{1},\dots,a_{n})$ to be the radical of the ideal $(a_{1},\dots,a_{n})$, and this is a concrete way to build the universal support.
 In particular, this shows that $D(a_{1},\dots,a_{n})=1$ if, and only if, $(a_{1},\dots,a_{n})=1$.
-It is also direct to see that $D(a,b)=D(a+b,ab)$ and more generally that
+It is also direct, for any support $d$, to see that $d(a,b)=d(a+b,ab)$ and more generally that
 $d(a_{1},\dots,a_{n})=d(s_{1},\dots,s_{n})\text{,}$
 where $s_{1},\dots,s_{n}$ are the elementary symmetric functions of $a_{1},\dots,a_{n}$.
-For the formulation of this theory of strictly Henselian rings, we need the notion of universal decomposition algebra $L=R[x_{1},\dots,x_{n}]=R[X_{1},\dots,X_{n}]/I$ of a monic polynomial
+For the formulation of this theory of strictly Henselian rings, we need the notion of universal splitting algebra [LQ, III, 4] $L=R[x_{1},\dots,x_{n}]=R[X_{1},\dots,X_{n}]/I$ of a monic polynomial
 $f=X^{n}-u_{1}X^{n-1}+u_{2}X^{n-2}$-$\dots$
 in $R[X]$, where $I$ is the ideal generated by $u_{1}-\sigma_{1},~\dots,~u_{n}-\sigma_{n}$, where $\sigma_{1},\dots,\sigma_{n}$ in $R[X_{1},\dots,X_{n}]$ are the elementary symmetric polynomials.
-We know that $L=R[x_{1},\dots,x_{n}]$ is freely generated as a module over $R$ and it follows from this that $L$ is a faithfully flat extension of $R$.
+We know, using Cauchy modules [LQ, III, 4.3], that $L=R[x_{1},\dots,x_{n}]$ is freely generated of rank $n!$ as a module over $R$ and it follows from this that $L$ is a faithfully flat extension of $R$.
+For the notion of faithfully flat extension, we follow the definition in [LQ, VIII, 6]: it is an $R$-algebra $i:R\rightarrow S$ such that
+  • if $u$ is a row vector in $A^{n}$ and $v$ a column vector in $B^{n}$ such that $i(u)v=0$ then we can find a matrix $M$ with coefficients in $A$ and $w$ such that $uA=0$ and $v=i(A)w$;
+  • if $i(a)$ belongs to the ideal $(i(a_{1}),\dots,i(a_{n}))$ in $B$ then $a$ belongs to the ideal $(a_{1},\dots,a_{n})$ in $A$.
+It then follows [LQ, VIII, 6] that $i$ is injective. It is also clear from this definition that $S$ is faithfully flat over $R$ when it is a finite free $R$-module of rank >0>0.
 We define $\delta_{i}(f)$ to be the element in $R$ equal to $\sigma_{i}(f^{\prime}(x_{1}),\dots,f^{\prime}(x_{n}))$ and $\Delta(f)$ to be $D(\delta_{1}(f),\dots,\delta_{n}(f))$.
@@
 Over a residually discrete local ring with residue field $F$, to be unramifiable means that the polynomial $f$ has a simple root in some extension of $F$.
+This holds in classical mathematics, but also constructively with a dynamical interpretation of extensions of $F$.
 Lemma 1.1.
-Let $R$ be a local ring residually discrete of maximal ideal $\mathsf{m}$ with residue field $F=R/\mathsf{m}$.
-If $f$ is a monic polynomial in $R[X]$ which has a simple root in some nontrivial $F$-algebra $A$, then $f$ is unramifiable.
+Let $R$ be a local ring, residually discrete with maximal ideal $\mathsf{m}$ and residue field $F=R/\mathsf{m}$.
+If $f$ is a monic polynomial in $R[X]$ which has a simple root in some nontrivial $F$-algebra AA, then $f$ is unramifiable.
 Lemma 1.2.
@@
 We write $h=(X-a_{1})\cdots(X-a_{p})$.
-We have in $R[x_{1},\dots,x_{n}]$, the decomposition algebra of $g$,
+We have in $R[x_{1},\dots,x_{n}]$, the universal splitting algebra of $g$,
 $1=D(f^{\prime}(a_{1}),\dots,f^{\prime}(a_{p}),h(x_{1})g^{\prime}(x_{1}),\dots,h(x_{n})g^{\prime}(x_{n}))$
@@
 $1=D(f^{\prime}(a_{1}),\dots,f^{\prime}(a_{p}))\vee\Delta(g)$
-in $R[x_{1},\dots,x_{n}]$. Since the decomposition algebra is faithfully flat, we get
+in $R[x_{1},\dots,x_{n}]$. Since the universal splitting algebra is faithfully flat, we get
 $1=D(f^{\prime}(a_{1}),\dots,f^{\prime}(a_{p}))\vee\Delta(g)$
@@
 If $a\neq u$ mod.
-$\mathsf{m}$ then we write $P=(X-a)Q$, and $u$ is residually a simple root of $Q$.
+$\mathsf{m}$ then we write P=(Xa)QP=(X-a)Q, and $u$ is residually a simple root of $Q$.
 ∎
@@
 Proof.
-Let $A=R[x_{1},\dots,x_{n}]$ be the universal decomposition algebra of $P$.
-This is a free $R$-module over $R$ of dimension $n!$ and the symmetric group acts on it [LQ, Fact 3-4.4].
+Let $A=R[x_{1},\dots,x_{n}]$ be the universal splitting algebra of $P$.
+This is a free $R$-module over $R$ of dimension $n!$ and the symmetric group acts on it [LQ, Fact III-4.4].
 We may identify $R[x]$ with the subalgebra $R[x_{1}]$ and we have $e(x_{1})$ in $R[x_{1}]$ which is idempotent mod.
@@
 The product $e(x_{1})\cdots e(x_{n})$ is in $R$ and it is an idempotent mod.
-$\mathsf{m}$, so it is $1$ or 00 mod.
+$\mathsf{m}$, so it is $1$ or 0 mod.
 $\mathsf{m}$.
@@
 Proof.
-Let $u$ be a vector column of generators of $M$, then we can find a vector $v$ in $N$ and a matrix $P$ with elements in $\mathsf{m}$ such that $u=v+Pu$.
-We then have $(I-P)u=v$ and hence the elements of $v$ are generators of $M$ since $\det(I-P)$ is invertible.
+Let $u$ be a vector column of generators of $M$: then we can find a vector $v$ in $N$ and a matrix $P$ with elements in $\mathsf{m}$ such that $u=v+Pu$.
+We then have (IP)u=v(I-P)u=v and hence the elements of $v$ are generators of $M$ since $\det(I-P)$ is invertible.
 ∎
 Corollary 2.1.
-Let $A$ be an $R$-algebra which is finitely generated as an $R$-module, and $u$ in $A$ which is invertible mod. $\mathsf{m}A$, then $u$ is invertible in $A$.
+Let $A$ be an $R$-algebra which is finitely generated as an $R$-module, and $u$ in $A$ which is invertible mod. $\mathsf{m}A$. Then $u$ is invertible in $A$.
 Proof.
@@
 We can find an idempotent $e$ in $R[x]$ mod.
-$\mathsf{m}R[x]$ such that $(f)=(e),~{}(g)=(1-e)$ mod.
+$\mathsf{m}R[x]$ such that (f)=(e),(g)=(1e)(f)=(e),~(g)=(1-e) mod.
 $\mathsf{m}R[x]$ [ALP, Prop.
@@
 In particular, $ex^{m}$ belongs to the $R$-module generated by $e,ex,\dots,ex^{m-1}$ (since $e$ is idempotent in $R[x]$) and we find in this way $G$ monic of degree $m$ such that $eG(x)=0$ in $R[x]$.
-Similarly we find $F$ monic of degree $n$ such that $(1-e)F(x)=0$ in $R[x]$.
+Similarly we find $F$ monic of degree $n$ such that (1e)F(x)=0(1-e)F(x)=0 in $R[x]$.
 We then have $F(x)G(x)=F(x)eG(x)=0$ and hence $P$ divides $FG$.
@@
 We decompose $F=X^{n}(1-X)^{m}h$ mod.
-$\mathsf{m}R[X]$, with $h$ monic and $h(0),~{}h(1)$ not in $\mathsf{m}$.
+$\mathsf{m}R[X]$, with $h$ monic and $h(0)$, $h(1)$ not in $\mathsf{m}$.
 If $n=0$ we have $F(0)=h(0)$ mod.
@@
 $\mathsf{m}A$, and hence invertible in $A$ using Corollary 2.1, and we can take $e=1$.
-Similarly, if $m=0$, we get that $1-a$ is invertible and we can take $e=0$.
+Similarly, if $m=0$, we get that 1a1-a is invertible and we can take $e=0$.
 In the remaining case, we can find using Lemma 2.3 a decomposition in monic polynomials $F=PQH$ in $R[X]$ with $P=X^{n},~Q=(1-X)^{m},~H=h$ mod.
 $\mathsf{m}$.
-We also have $U$ and $V$ in $R[X]$ such that $X(1-X)U+HV=Res(X(1-X),H)=1$ [LQ, Fact 3-7.1], [ALP, Lemma 3.10], and so $H(a)V(a)=1$ mod.
+We also have $U$ and $V$ in $R[X]$ such that X(1X)U+HV=Res(X(1X),H)=1X(1-X)U+HV=Res(X(1-X),H)=1 [LQ, Fact 3-7.1], [ALP, Lemma 3.10], and so $H(a)V(a)=1$ mod.
 $\mathsf{m}A$, and so $H(a)$ is invertible mod.
@@
 Since $F(a)=P(a)Q(a)H(a)=0$, we have $P(a)Q(a)=0$ in $A$.
-We also have $P(a)=a$ and $Q(a)=1-a$ mod.
+We also have $P(a)=a$ and Q(a)=1aQ(a)=1-a mod.
 $\mathsf{m}A$ and so $P(a)+Q(a)=1$ mod.
@@
   1. 1.
-    $A$ is projective of finite type;
+    AA is projective of finite type;
   2. 2.
     the canonical map $A\otimes_{R}A^{op}\rightarrow End_{k}(A)$ is an isomorphism.
-Note that we allow the trivial algebra to be an Azumaya algebra.
+Note that we allow the trivial algebra to be an Azumaya algebra. Note also that any matrix algebra is Azumaya.
+One can equivalently describe the algebra structure by a bilinear map $\mu:A\times A\rightarrow A$; its scalar extension along $k\rightarrow S$ is the unique $S$-algebra $A_{S}$ through which every bilinear map extending $\mu$ factors.
+Thus all subsequent arguments can be phrased without explicit tensor notation.
+Over a finite free $k$-module, to give an algebra structure is represented concretely by a multiplication table.
+A module $A$, projective of finite type over $k$, is given concretely by a projection matrix over $k$ [LQ], and we can similarly represent the algebra structure by a multiplication table (so $m^{3}$ elements of $k$ if the projection matrix has finite size $m$).
+We can then compute [LQ, X, Theorem 1.6] comaximal elements $s_{1},\dots,s_{l}$ such that $A_{s_{i}}$ is free over $k_{s_{i}}$, and the second condition for being an Azumaya algebra becomes then that some determinants $\delta_{1},\dots,\delta_{l}$ are invertible [^5] After localisation to $s_{j}$ the Azumaya algebra becomes free of rank $r_{i}=n_{i}^{2}\leqslant m$..
+From this discussion follows that if $A$ is an Azumaya algebra, given by a multiplication table, and we have a $k$-algebra $k\rightarrow S$ then the algebra $A_{S}$ over $S$, with the same multiplication in $S$, is also an Azumaya algebra.
 We have the following characterisation, which generalizes the result that an algebra is central simple over a field if, and only if, it can be split by a separable extension of this field.
 Theorem 3.1.
-$A$ is an Azumaya algebra over $k$ if, and only if, we can build a finite tree of root $k$ using the following, where all leaves are matrix algebras:
+AA is an Azumaya algebra over $k$ if, and only if, we can build a finite tree with root $k$ using the following, where at all leaves all scalar extensions $A_{S}$ are matrix algebras:
   1. 1.
@@
     $S$ is covered by $S\rightarrow S[X]/(P)$ if $P$ is monic and unramifiable.
+We prove first that if we have such a tree then the algebra $A$ is Azumaya.
 If we have such a tree starting from $k$, notice that each extension $S\rightarrow S[X]/(P)$ for $P$ monic is faithfully flat, and $S\rightarrow\prod_{i}S[1/u_{i}]$ if $1=(u_{1},\dots,u_{n})$ is also faithfully flat.
@@
 VIII-6.7].
-The map $A\otimes_{k}A^{op}\rightarrow End_{k}(A)$ is then a bijection since it becomes a bijection by faithfully flat extension.
-The converse follows from the following two results.
+The map $A\otimes_{k}A^{op}\rightarrow End_{k}(A)$ is then a bijection since it becomes a bijection by a faithfully flat extension.
+(As we saw before, to be a bijection is witnessed by the fact that some determinants are invertible, and an element is invertible if it becomes invertible by a faithfully flat extension.)
+For the converse, we look at the algebra $A$ (given concretely by its multiplication table) inside the étale topos over $k$, using the site defined in the first section.
+(This is similar to the constructive use of the algebraic closure in [MC].) The generic ring $R$, defined by $R(S)=S$ for $S$ a finitely presented $k$-algebra, is, in the internal logic of this topos, a local strictly Henselian ring.
+The algebra $A$ defines then an $R$-algebra $A_{R}(S)=A_{S}$, which is an Azumaya algebra, given by the same multiplication table as the one of $A$.
+If we can prove in intuitionistic logic that an Azumaya algebra over a strictly Henselian ring is a matrix algebra, we will get, using the interpretation of intuitionistic logic in a sheaf model [Joh, MC], a covering tree of $k$ which gives Theorem 3.1.
+Internally in the classifying topos of stricly Henselian $k$-algebras, the object $R$ is local strictly Henselian and the claim that $A_{R}$ is a matrix algebra is a coherent statement, provable intuitionistically.
+Such a proof yields externally a finite covering tree of $k$ satisfying the conditions above.
+The converse is a consequence of the following two results.
 Lemma 3.1.
@@
 We have that $Ae$ is projective and hence free since $R$ is a local ring.
-The map $A\rightarrow End_{R}(Ae)$ is then an isomorphism, since it is a map between two free $R$-modules of the same dimension, and it is residually an isomorphism.
+The map $A\rightarrow End_{R}(Ae)$ is then an isomorphism, since it is a map between two free $R$-modules of same dimension, and it is residually an isomorphism.
 ∎
+The next proof uses Barr’s Theorem, similar to the way it is used in [Penon, Wraith], but the difference is that we only need a constructively valid version of this theorem.
 Corollary 3.1.
@@
 Proof.
-For a given Azumaya algebra, the conclusion is coherent, and the hypothesis that $A$ is an Azumaya algebra is formulated in a coherent theory.
-Since it is proved with the extra hypothesis that $R$ is residually discrete, the syntactic version of Barr’s Theorem [^3] We can actually follow the argument in [Joh] and apply Friedman-Dragalin’s translation [Avi] w.r.t.
-to the proposition expressing that $A$ is a matrix algebra.
+For a given Azumaya algebra, the conclusion is coherent [Joh, MC], and the hypothesis that $A$ is an Azumaya algebra is formulated in a coherent theory.
+Since it is proved with the extra hypothesis that $R$ is residually discrete, the syntactic version of Barr’s Theorem [^6] We can actually follow the argument in [Joh] and apply Friedman-Dragalin’s translation [Avi] with respect to the proposition expressing that $A$ is a matrix algebra.
 (constructively valid) shows that it can be proved without this hypothesis.
 ∎
-Theorem 3.1 follows by interpreting this result in the classifying topos of the theory of strictly Henselian $k$-algebra.
+Theorem 3.1 follows by interpreting this result in the classifying topos of the theory of strictly Henselian $k$-algebras.
 The statement of Theorem 3.1 can be seen as a generalisation of the fact that over a field, a central simple algebra becomes a matrix algebra by a faithfully flat extension [CLN] (classically over an algebraic extension).
@@
 Lemma 3.3.
-If $R$ is a local ring and $A$ a matrix algebra of rank $>0$ [LQ, Def. X-2.2] then any automorphism of $A$ is of the form $x\mapsto axa^{-1}$ for some $a$ in $GL_{n}(A)$.
+If $R$ is a local ring and $A$ a matrix algebra of rank $\geqslant 1$ [LQ, Def. X-2.2] then any automorphism of $A$ is of the form $x\mapsto axa^{-1}$ for some $a$ in $GL_{n}(A)$.
 Proof.
@@
 ∎
+For the next result, we use the local-global principle and the descent principle with respect to faithfully flat extension for projective modules ([LQ, Theorem VIII-6.7]).
+For completeness, we recall the statement.
+Note that the proof in [LQ] is constructive, using the concrete definition of faithfully flat extension and the fact that, also constructively, a module is finite projective if, and only if, it is finitely presented and flat.
+Theorem 3.2 (Descent).
+Let $A\rightarrow B$ be a faithfully flat extension and $M$ an $A$-module.
+If $M\otimes_{A}B$ is finite projective, then so is $M$.
+If we have $1=(s_{1},\dots,s_{n})$ in $A$ and each $M_{s_{i}}$ is finite projective over $A_{s_{i}}$ then so is $M$.
 Theorem 3.3 (Skolem-Noether).
-If $A$ is an Azumaya algebra, of rank $>0$ as a projective module over a ring $k$, and $\psi$ is an automorphism of $A$, then $M=\{a\in A~{}|~{}\forall_{x}ax=\psi(x)a\}$ is a projective $k$-module of rank $1$.
+If $A$ is an Azumaya algebra, of rank $\geqslant 1$ as a projective module over a ring $k$, and $\psi$ is an automorphism of $A$, then $M=\{a\in A~|~\forall_{x}\,ax=\psi(x)a\}$ is a projective $k$-module of rank $1$.
 Proof.
-We use Theorem 3.1.
-The statement holds on the leaves ($M$ is even free of rank $1$ on the leaves by Lemma 3.2) and descends to the root by tree induction using the local-global principle and the descent principle w.r.t.
-faithfully flat extension for projective modules ([LQ, Thm.
-VIII-6.7]).
-∎
+We use Theorem 3.1. The statement holds on the leaves ($M$ is even free of rank 11 on the leaves by Lemma 3.2) and descends to the root by tree induction using Theorem 3.2. ∎
 If $k$ is local, the module $M$ is free of rank $1$ and any generator $u$ of $M$ is invertible (since it is invertible on the leaves) and $\psi$ is the inner automorphism defined by $u$, so we recover the usual form of the Skolem-Noether Theorem (any automorphism of a central simple algebra is an inner automorphism).
-Claude Quitté suggested the following refinement.
+Claude Quitté has suggested the following refinement.
 Theorem 3.4.
-If $A$ is an Azumaya algebra, of rank $>0$ as a projective module over a ring $k$, and $\psi$ is an automorphism of $A$, then
+If $A$ is an Azumaya algebra, of rank $\geqslant 1$ as a projective module over a ring $k$, and $\psi$ is an automorphism of $A$, then
 $M=\{a\in A~|~\forall_{x}\,ax=\psi(x)a\}$ and $N=\{a\in A~|~\forall_{x}\,a\psi(x)=xa\}$
@@
 Proof.
-We have on the leaves $MN\subseteq k1$ and this statement descends to the root by tree induction.
+We have on the leaves $M_{S}N_{S}\subseteq S1$ and this statement descends to the root by tree induction.
 This can be used to define a map $M\rightarrow Hom(N,k),~x\mapsto(y\mapsto xy)$.
@@
 Theorem 4.1.
-$A$ is an Azumaya algebra over $k$ if, and only if, we can build a finite tree of root $k$ using the following, where all leaves are matrix algebras:
+AA is an Azumaya algebra over $k$ if, and only if, we can build a finite tree of root $k$ using the following, where all leaves are matrix algebras:
   1. 1.

### v4

@@
-Azumaya algebras and Barr’s Theorem
+arXiv:2306.17679v4[math.AC] 01 Mar 2026
+arXiv:2306.17679v4[math.AC] 01 Mar 2026
 Azumaya algebras and Barr’s Theorem
 Thierry Coquand Henri Lombardi Stefan Neuwirth
-(November 24, 2025)
 Introduction
-The goal of this paper is to show how one can use the notion of classifying topos of a coherent theory [Joh, CLR, Wraith], here the theory of strictly Henselian local rings [Wraith], to prove some basic results about Azumaya algebras in a constructive setting.
-A fundamental invariant associated to any commutative field $F$ is its Brauer group, whose elements are division algebras over $F$, up to isomorphism.
+The goal of this paper is to show how one can use the notion of classifying topos of a coherent theory [Joh, CLR, Wraith], here the theory of separably closed local rings [Wraith], to prove some basic results about Azumaya algebras in a constructive setting.
+A fundamental invariant associated to a commutative field $F$ is its Brauer group, whose elements are the division algebras over $F$, up to isomorphism.
 Alternatively an element of the Brauer group can be described as a central simple algebra, up to Morita equivalence, and the group operation is then the tensor product.
-Two important characterisations of the notion of central simple algebra are the following [^1] We study these characterisations in a constructive setting in the reference [CLN].:
+Two important characterisations of the notion of central simple algebra are the following: [^1] We study these characterisations in a constructive setting in the reference [CLN].
   1. 1.
-    an algebra $A$ over $F$ which is finite and such that the canonical map $A\otimes_{F}A^{op}\rightarrow End_{F}(A)$ is an isomorphism; or, equivalently,
+    an algebra $A$ over $F$ which is finitely generated and such that the canonical map $A\otimes_{F}A^{\mathrm{op}}\rightarrow\operatorname{End}_{F}(A)$ is an isomorphism; or, equivalently,
   2. 2.
     an algebra $A$ which becomes a matrix algebra over an algebraic extension of $F$, or, equivalently, which becomes a matrix algebra over an algebraic and separable extension of $F$.
-This notion of central simple algebra over a field has been generalized to the notion of Azumaya algebra over an arbitrary commutative ring $R$.
-The definition is a generalisation of the first characterisation: an algebra $A$ which is projective of finite type over $R$ and such that the canonical map $A\otimes_{F}A^{op}\rightarrow End_{R}(A)$ is an isomorphism.
-In this note, we show in a constructive setting that this definition is equivalent to a suitable generalisation of the second characterisation, using a constructively valid version of Barr’s Theorem [Joh].
+This notion of central simple algebra over a field has been generalised to the notion of Azumaya algebra over an arbitrary commutative ring $k$.
+The definition is a generalisation of the first characterisation: an algebra $A$ which is finitely generated projective over the ring $k$ and such that the canonical map $A\otimes_{k}A^{\mathrm{op}}\rightarrow\operatorname{End}_{k}(A)$ is an isomorphism.
+In this paper, we show in a constructive setting that this definition is equivalent to a suitable generalisation of the second characterisation, using a constructively valid version of Barr’s Theorem [Joh].
 1 Some coherent theories
-We will consider some coherent theories [Joh] in the language of rings.
+We consider some coherent theories [Joh] in the language of rings.
 1.1 Theory of local rings
 The theory of local rings can be formulated in a coherent way with the axioms
-$0\neq 1~~~~~~~~~\mathsf{inv}(x)\vee\mathsf{inv}(1-x)$
+$0\neq 1\qquad\qquad\mathsf{inv}(x)\vee\mathsf{inv}(1-x)$
 where $\mathsf{inv}(x)$ denotes $\exists_{y}~1=xy$.
-If we add to our language a new predicate $\mathsf{m}(x)$ with the axioms
-$\mathsf{inv}(x)\vee\mathsf{m}(x)~~~~~~~~~~\neg(\mathsf{m}(x)\wedge\mathsf{inv}(x))$
-then we get the theory of local rings that are residually discrete [^2] The terminology “discrete” for a set with a decidable equality, while conflicting with some topological intuitions, is well established in constructive mathematics since the work of Bishop [Bish], and we follow this usage in this paper., i.e.
-where the property of being invertible is now decidable, and the predicate $\mathsf{m}(x)$ corresponds to the unique maximal ideal of the ring.
-In this theory, we can express in a coherent way that the ring is Henselian by stating that for any monic polynomial $P$, if we have $u$ such that $P(u)\in\mathsf{m}$ and $P^{\prime}(u)\notin\mathsf{m}$, we can find $\alpha$ such that $P(\alpha)=0$ and $\alpha-u\in\mathsf{m}$.
+If we add to our language a new predicate $\mathfrak{m}(x)$ with axioms
+$\mathsf{inv}(x)\vee\mathfrak{m}(x)\qquad\qquad\neg(\mathfrak{m}(x)\wedge\mathsf{inv}(x))$
+then we get the theory of local rings that are residually discrete, [^2] The terminology “discrete” for a set with a decidable equality, while conflicting with some topological intuitions, is well established in constructive mathematics since the work of Bishop [Bish], and we follow this usage in this paper.
+i.e.
+where the property of being invertible is now decidable, and the predicate $\mathfrak{m}(x)$ corresponds to the unique maximal ideal of the ring.
+In this theory, we can express in a coherent way that the ring is Henselian by stating that for any monic polynomial $f$, if we have $u$ such that $f(u)\in\mathfrak{m}$ and $f^{\prime}(u)\notin\mathfrak{m}$ then we can find $\alpha$ such that $f(\alpha)=0$ and $\alpha-u\in\mathfrak{m}$.
 Equivalently [ALP], any monic polynomial of degree $l+1$ which is of the form $X^{l}(X-1)$ mod.
-$\mathsf{m}$ has a root which is equal to $1$ mod.
-$\mathsf{m}$.
-The work [ALP] shows in an intuitionistic setting for this theory the lifting property for idempotents of a finitely generated algebra over a Henselian local ring.
-It does not seem possible however to express the property of being Henselian in a coherent way in the language of rings (i.e.
-without introducing the predicate $\mathsf{m}(x)$).
-Surprisingly, as noticed by G.
-Wraith [Wraith], the property of being strictly Henselian, corresponding classically to the property of being Henselian with the residue field being separably closed, can be expressed in a coherent way.
-In this note, we present some remarks connected to this observation.
+$\mathfrak{m}$ has a root which is equal to $1$ mod.
+$\mathfrak{m}$.
+The work [ALP] shows in a constructive setting for this theory the lifting property for idempotents of a finitely generated commutative algebra over a Henselian local ring.
+It does not seem possible however to express the property of being Henselian in a coherent way in the language of local rings (i.e.
+without introducing the predicate $\mathfrak{m}(x)$).
+Surprisingly, as observed by G.
+Wraith [Wraith], the property of being separably closed, corresponding classically to the property of being Henselian with the residue field being separably closed, can be expressed in a coherent way.
+In this paper, we present some remarks connected to this observation.
 The first one is a small variation of Wraith’s axiomatisation.
 Both axiomatisations can be used to define the étale site over a ring.
-Finally, we show in a constructive setting that Azumaya algebras over a ring are exactly the algebras that are locally matrix algebras for the étale topology.
+Finally, we show in a constructive setting that Azumaya algebras are exactly the algebras that are locally matrix algebras for the étale topology.
 1.2 Theory of separably closed local rings
 If $R$ is a ring, and $a_{1},\dots,a_{n}$ elements of $R$, we write $(a_{1},\dots,a_{n})$ for the ideal of $R$ generated by $a_{1},\dots,a_{n}$.
-We define, following Joyal [Joy], the Zariski spectrum of $R$ as the universal support $D:R\rightarrow Z(R)$ over $R$.
-A support is a map $d:R\rightarrow M$ into a bounded distributive lattice $M$ satisfying the relations
-$d(1)=1~~~~~d(0)=0~~~~~~d(a+b)\leqslant d(a)\vee d(b)~~~~~~~~d(ab)=d(a)\wedge d(b)$
-We write $d(a_{1},\dots,a_{n})$ for $d(a_{1})\vee\dots\vee d(a_{n})$.
+We define, following Joyal [Joy], the Zariski spectrum of $R$ as the universal support $D\colon R\rightarrow Z(R)$ over $R$.
+A support is a map $d\colon R\rightarrow M$ into a bounded distributive lattice $M$ satisfying the relations
+$d(1)=1\qquad\qquad d(0)=0\qquad\qquad d(a+b)\leqslant d(a)\vee d(b)\qquad\qquad d(ab)=d(a)\wedge d(b)$
+We write $d(a_{1},\dots,a_{n})$ for $d(a_{1})\vee\cdots\vee d(a_{n})$.
 It is direct that $d(a)\leqslant d(a_{1},\dots,a_{n})$ as soon as $a$ belongs to the radical of the ideal $(a_{1},\dots,a_{n})$.
@@
 It is also direct, for any support $d$, to see that $d(a,b)=d(a+b,ab)$ and more generally that
-$d(a_{1},\dots,a_{n})=d(s_{1},\dots,s_{n})\text{,}$
-where $s_{1},\dots,s_{n}$ are the elementary symmetric functions of $a_{1},\dots,a_{n}$.
-For the formulation of this theory of strictly Henselian rings, we need the notion of universal splitting algebra [LQ, III, 4] $L=R[x_{1},\dots,x_{n}]=R[X_{1},\dots,X_{n}]/I$ of a monic polynomial
-$f=X^{n}-u_{1}X^{n-1}+u_{2}X^{n-2}$-$\dots$
-in $R[X]$, where $I$ is the ideal generated by $u_{1}-\sigma_{1},~\dots,~u_{n}-\sigma_{n}$, where $\sigma_{1},\dots,\sigma_{n}$ in $R[X_{1},\dots,X_{n}]$ are the elementary symmetric polynomials.
-We know, using Cauchy modules [LQ, III, 4.3], that $L=R[x_{1},\dots,x_{n}]$ is freely generated of rank $n!$ as a module over $R$ and it follows from this that $L$ is a faithfully flat extension of $R$.
-For the notion of faithfully flat extension, we follow the definition in [LQ, VIII, 6]: it is an $R$-algebra $i:R\rightarrow S$ such that
-  • if $u$ is a row vector in $A^{n}$ and $v$ a column vector in $B^{n}$ such that $i(u)v=0$ then we can find a matrix $M$ with coefficients in $A$ and $w$ such that $uA=0$ and $v=i(A)w$;
-  • if $i(a)$ belongs to the ideal $(i(a_{1}),\dots,i(a_{n}))$ in $B$ then $a$ belongs to the ideal $(a_{1},\dots,a_{n})$ in $A$.
-It then follows [LQ, VIII, 6] that $i$ is injective. It is also clear from this definition that $S$ is faithfully flat over $R$ when it is a finite free $R$-module of rank >0>0.
+$d(a_{1},\dots,a_{n})=d(c_{1},\dots,c_{n})\text{,}$
+where $c_{1},\dots,c_{n}$ are the elementary symmetric functions of $a_{1},\dots,a_{n}$.
+For the formulation of the theory of separably closed local rings, we need the notion of universal splitting algebra [LQ, Sect.
+III-4] $L=R[x_{1},\dots,x_{n}]=R[X_{1},\dots,X_{n}]/I$ of a monic polynomial
+$f=X^{n}-u_{1}X^{n-1}+u_{2}X^{n-2}$-$\cdots$
+in $R[X]$, where $I$ is the ideal generated by $u_{1}-\sigma_{1}$, …, $u_{n}-\sigma_{n}$, with $\sigma_{1}$, …, $\sigma_{n}$ in $R[X_{1},\dots,X_{n}]$ the elementary symmetric polynomials.
+We know, using Cauchy modules [LQ, Fact III-4.3], that $L$ is freely generated of rank $n!$ as a module over $R$ and it follows from this that $L$ is a faithfully flat extension of $R$.
+For the notion of faithfully flat extension, we follow the definition in [LQ, Sect. VIII-6]: it is a commutative $R$-algebra $i\colon R\rightarrow S$ such that
+  • if $u$ is a row vector in $R^{n}$ and $v$ a column vector in $S^{n}$ such that $i(u)v=0$ then we can find a matrix $M$ with coefficients in $R$ and $w$ such that $uM=0$ and $v=i(M)w$;
+  • if $i(a)$ belongs to the ideal $(i(a_{1}),\dots,i(a_{n}))$ in $S$ then $a$ belongs to the ideal $(a_{1},\dots,a_{n})$ in $R$.
+It then follows [LQ, Sect. VIII-6] that $i$ is injective. It is also clear from this definition that $S$ is faithfully flat over $R$ when it is a finitely generated free $R$-module of rank >0>0.
 We define $\delta_{i}(f)$ to be the element in $R$ equal to $\sigma_{i}(f^{\prime}(x_{1}),\dots,f^{\prime}(x_{n}))$ and $\Delta(f)$ to be $D(\delta_{1}(f),\dots,\delta_{n}(f))$.
 We say that $f$ is unramifiable if, and only if, $1=\Delta(f)$.
-Since $D(f^{\prime}(x_{1}),\dots,f^{\prime}(x_{n}))=D(\delta_{1}(f),\dots,\delta_{n}(f))$, this condition is equivalent to $(f^{\prime}(x_{1}),\dots,f^{\prime}(x_{n}))=1$ in $L$.
-Over a residually discrete local ring with residue field $F$, to be unramifiable means that the polynomial $f$ has a simple root in some extension of $F$.
-This holds in classical mathematics, but also constructively with a dynamical interpretation of extensions of $F$.
+Since $\Delta(f)=D(f^{\prime}(x_{1}),\dots,f^{\prime}(x_{n}))$, this condition is equivalent to $(f^{\prime}(x_{1}),\dots,f^{\prime}(x_{n}))=1$ in the universal splitting algebra $L$.
+Over a residually discrete local ring with maximal ideal $\mathfrak{m}$, to be unramifiable means that the polynomial $f$ has a simple root in some extension of the residue field $R/\mathfrak{m}$.
+This holds in classical mathematics, but also constructively with a dynamical interpretation of extensions of $R/\mathfrak{m}$.
 Lemma 1.1.
-Let $R$ be a local ring, residually discrete with maximal ideal $\mathsf{m}$ and residue field $F=R/\mathsf{m}$.
-If $f$ is a monic polynomial in $R[X]$ which has a simple root in some nontrivial $F$-algebra AA, then $f$ is unramifiable.
-Lemma 1.2.
-If $f=(X-a_{1})\dots(X-a_{p})g$ in $R[X]$ with $g$ monic of degree $n$ and $\Delta(f)=1$ and $R$ is local, then $f^{\prime}(a_{1})$ is invertible or $\dots$ or $f^{\prime}(a_{p})$ is invertible or $\Delta(g)=1$.
+Let $R$ be a local ring, residually discrete with maximal ideal $\mathfrak{m}$ and residue field $R/\mathfrak{m}$.
+If $g$ is a monic polynomial in $R[X]$ which has a simple root in some nontrivial commutative $R/\mathfrak{m}$-algebra, then $g$ is unramifiable.
+Lemma 1.2 ([LQ2, Lemme IX-7.1]).
+If $g=(X-a_{1})\cdots(X-a_{p})f$ in $R[X]$ with $f$ monic of degree $n$, $\Delta(g)=1$, and $R$ is local, then $g^{\prime}(a_{1})$ is invertible or … or $g^{\prime}(a_{p})$ is invertible or $\Delta(f)=1$.
 Proof.
 We write $h=(X-a_{1})\cdots(X-a_{p})$.
-We have in $R[x_{1},\dots,x_{n}]$, the universal splitting algebra of $g$,
-$1=D(f^{\prime}(a_{1}),\dots,f^{\prime}(a_{p}),h(x_{1})g^{\prime}(x_{1}),\dots,h(x_{n})g^{\prime}(x_{n}))$
+We have in $R[x_{1},\dots,x_{n}]$, the universal splitting algebra of $f$,
+$1=D(g^{\prime}(a_{1}),\dots,g^{\prime}(a_{p}),h(x_{1})f^{\prime}(x_{1}),\dots,h(x_{n})f^{\prime}(x_{n}))$
 and hence
-$1=D(f^{\prime}(a_{1}),\dots,f^{\prime}(a_{p}),g^{\prime}(x_{1}),\dots,g^{\prime}(x_{n}))$
+$1=D(g^{\prime}(a_{1}),\dots,g^{\prime}(a_{p}),f^{\prime}(x_{1}),\dots,f^{\prime}(x_{n}))$
 and from this follows
-$1=D(f^{\prime}(a_{1}),\dots,f^{\prime}(a_{p}))\vee\Delta(g)$
+$1=D(g^{\prime}(a_{1}),\dots,g^{\prime}(a_{p}))\vee\Delta(f)$
 in $R[x_{1},\dots,x_{n}]$. Since the universal splitting algebra is faithfully flat, we get
-$1=D(f^{\prime}(a_{1}),\dots,f^{\prime}(a_{p}))\vee\Delta(g)$
+$1=D(g^{\prime}(a_{1}),\dots,g^{\prime}(a_{p}))\vee\Delta(f)$
 in $R$, and since $R$ is local we have the conclusion. ∎
-Lemma 1.3.
-If $R$ is local and any monic unramifiable polynomial has a root in $R$ then any monic unramifiable polynomial has a simple root in $R$.
+Definition 1.1 ([LQ2, Définition IX-7.2]).
+A local ring $R$ is separably closed if every unramifiable monic polynomial has a root in $R$.
+Lemma 1.3 ([LQ2, Lemme IX-7.3]).
+If $R$ is a separably closed local ring then any unramifiable monic polynomial has a simple root in $R$.
 Proof.
-Assume that $R$ is local and any monic unramifiable polynomial has a root in $R$.
-Let $f$ be a monic unramifiable polynomial.
-It has a root $a_{1}$ and we can write $f=(X-a_{1})f_{1}$.
-By Lemma 1.2, $a_{1}$ is a simple root of $f$ or $\Delta(f_{1})=1$.
-If $\Delta(f_{1})=1$ then it has a root $a_{2}$ and $f=(X-a_{1})(X-a_{2})f_{2}$.
-Then by Lemma 1.2 again, $a_{1}$ or $a_{2}$ is a simple root of $f$ or $\Delta(f_{2})=1$, and so on until we find a simple root of $f$.
+Assume that $R$ is separably closed local.
+Let $g$ be a monic unramifiable polynomial.
+It has a root $a_{1}$ and we can write $g=(X-a_{1})f_{1}$.
+By Lemma 1.2, $a_{1}$ is a simple root of $g$ or $\Delta(f_{1})=1$.
+If $\Delta(f_{1})=1$ then it has a root $a_{2}$ and $g=(X-a_{1})(X-a_{2})f_{2}$.
+Then by Lemma 1.2 again, $a_{1}$ or $a_{2}$ is a simple root of $g$ or $\Delta(f_{2})=1$, and so on until we find a simple root of $g$.
 ∎
-Let $k$ be a ring. We consider now the category of finitely presented $k$-algebras with the covering
+Let $k$ be a ring. We can consider the following two coherent theories for a commutative $k$-algebra $R$:
   1. 1.
-    $S$ is covered by $S\rightarrow S[1/u_{1}],\dots,S\rightarrow S[1/u_{m}]$ if $1=(u_{1},\dots,u_{m})$;
+    $T_{1}$ states that $R$ is local, and that any monic unramifiable polynomial has a root.
   2. 2.
-    $S$ is covered by $S\rightarrow S[X]/(P)$ if $P$ is monic and unramifiable.
-It also follows from Lemma 1.3 that we can instead take the following covering, since it classifies an equivalent theory.
+    $T_{2}$ states that $R$ is local, and that any monic unramifiable polynomial has a simple root.
+We have just seen that these two theories are equivalent.
+The equivalence between these two theories can be seen as a logical formulation of the “trick” attributed to O. Gabber described in [Stacks, Tag 02LI].
+A morphism
+$S\longrightarrow S[X]/(f),$
+with $f$ monic, is finite locally free, and thus represents the “finite locally free” stage in Gabber’s refinement of an étale covering.
+Étaleness is obtained by further localising, that is, by passing to the Zariski covering
+$S\longrightarrow S[X]/(f)[1/f^{\prime}].$
+Lemma 1.3 provides the logical ingredient of this refinement: over a local base ring, the existence of a root of a monic unramifiable polynomial implies the existence of a simple root.
+Internally, this shows that coverings generated by finite locally free morphisms of the above form can be refined, after a Zariski localisation, to étale morphisms.
+In this sense, the equivalence between the two theories $T_{1}$ and $T_{2}$ is the internal counterpart of Gabber’s trick: finite locally free coverings suffice to generate the same local objects as étale coverings, up to Zariski refinement.
+These two coherent theories define two different sites (notions of covering), in the sense of Grothendieck (see e.g.
+[Joh]), over the opposite of the category of finitely presented commutative $k$-algebras [Joh].
+The first theory $T_{1}$ corresponds to:
   1. 1.
-    $S$ is covered by $S\rightarrow S[1/u_{1}],\dots,S\rightarrow S[1/u_{m}]$ if $1=(u_{1},\dots,u_{m})$;
+    $S$ is covered by $\rightarrow S[1/s_{1}], …, SS$$\rightarrow S[1/s_{n}$] if $1=(s_{1},\dots,s_{n})$;
   2. 2.
-    $S$ is covered by $S\rightarrow S[X]/(P)[1/P^{\prime}]$ if $P$ is monic and unramifiable.
-This defines a site, in the sense of Grothendieck (see e.g.
-[Joh]), over the opposite of the category of finitely presented $k$-algebras.
-The corresponding sheaf model is the classifying topos of the theory of strictly Henselian $k$-algebras [Joh].
-In this sheaf model, we have a “generic” $k$-algebra which is strictly Henselian.
+    $S$ is covered by $S\rightarrow S[X]/(f)$ if $f$ is monic and unramifiable.
+The second theory $T_{2}$ corresponds to:
+  1. 1.
+    $S$ is covered by $\rightarrow S[1/s_{1}], …, SS$$\rightarrow S[1/s_{n}$] if $1=(s_{1},\dots,s_{n})$;
+  2. 2.
+    $S$ is covered by $S\rightarrow S[X]/(f)[1/f^{\prime}]$ if $f$ is monic and unramifiable.
+Both define the same corresponding classifying topos of the theory of separably closed local commutative $k$-algebras [Joh].
+In this sheaf model, we have a “generic” local commutative $k$-algebra $R$ which is separably closed.
 We also may state, following Wraith,
-Theorem 1.1.
-If $R$ is a local ring which is also residually discrete, then $R$ is strictly Henselian if, and only if, $R$ is Henselian and its residue field is separably closed.
+Theorem 1.1 ([LQ2, Lemmes IX-7.4, IX-7.5]).
+If $R$ is a local ring which is also residually discrete, then $R$ is separably closed if, and only if, $R$ is Henselian and its residue field is separably closed.
 Proof.
-Let $P$ be a monic polynomial which has residually a simple root $u$.
-We show by induction on the degree of $P$ that $P$ has a root in $R$ which is equal to $u$ mod.
-$\mathsf{m}$.
-The polynomial $P$ is unramifiable by Lemma 1.1, since it has residually a simple root, and it has a root $a$.
+Let $g$ be a monic polynomial which has residually a simple root $u$.
+We show by induction on the degree of $g$ that $g$ has a root in $R$ which is equal to $u$ mod.
+$\mathfrak{m}$.
+The polynomial $g$ is unramifiable by Lemma 1.1, since it has residually a simple root, and it has a root $a$.
 If $a\neq u$ mod.
-$\mathsf{m}$ then we write P=(Xa)QP=(X-a)Q, and $u$ is residually a simple root of $Q$.
+$\mathfrak{m}$ then we write g=(Xa)fg=(X-a)f, and $u$ is residually a simple root of $f$.
 ∎
+Thus a residually discrete local ring is separably closed if and only if it is “strictly Henselian”.
 2 Some lemmas about Henselian rings
-In the next section, we need some results about Henselian rings.
-In order to make the paper self-contained we list here these results with their proofs.
-They are taken from the reference [ALP], with a variation in the proof of Lemma 2.1.
-We assume in this section that $R$ is local residually discrete with maximal ideal $\mathsf{m}$ and Henselian.
-Lemma 2.1 ([ALP, Prop. 5.1]).
-Let $P$ be a monic polynomial of degree $n$ in $R[X]$ and $R[x]=R[X]/(P)$.
+In the next section, we need a generalisation of a result about Henselian rings. In order to make the paper self-contained we mention here this result, Theorem 2.1, taken from the reference [ALP].
+We assume in this section that $R$ is local residually discrete with maximal ideal $\mathfrak{m}$ and Henselian.
+Lemma 2.1 is a main step in the proof of Theorem 2.1: we provide a variation of the proof that requires less constructive Galois theory than the original one in [ALP].
+Lemma 2.1 ([ALP, Prop. 5.7]).
+Let $f$ be a monic polynomial of degree $n$ in $R[X]$ and $R[x]=R[X]/(f)$.
 If we have $e(X)$ in $R[X]$ such that $e(x)$ is idempotent mod.
-$\mathsf{m}R[x]$, then we can find $u$ idempotent in $R[x]$ and $u=e(x)$ mod.
-$\mathsf{m}R[x]$.
+$\mathfrak{m}R[x]$, then we can find $u$ idempotent in $R[x]$ with $u=e(x)$ mod.
+$\mathfrak{m}R[x]$.
 Proof.
-Let $A=R[x_{1},\dots,x_{n}]$ be the universal splitting algebra of $P$.
+Let $L=R[x_{1},\dots,x_{n}]$ be the universal splitting algebra of $f$.
 This is a free $R$-module over $R$ of dimension $n!$ and the symmetric group acts on it [LQ, Fact III-4.4].
 We may identify $R[x]$ with the subalgebra $R[x_{1}]$ and we have $e(x_{1})$ in $R[x_{1}]$ which is idempotent mod.
-$\mathsf{m}R[x_{1}]$.
+$\mathfrak{m}R[x_{1}]$.
 We show that we have $u_{1}$ in $R[x_{1}]$ which is idempotent in $R[x_{1}]$ and equal to $e(x_{1})$ mod.
-$\mathsf{m}R[x_{1}]$.
+$\mathfrak{m}R[x_{1}]$.
 The product $e(x_{1})\cdots e(x_{n})$ is in $R$ and it is an idempotent mod.
-$\mathsf{m}$, so it is $1$ or 0 mod.
-$\mathsf{m}$.
+$\mathfrak{m}$, so it is $1$ or 0 mod.
+$\mathfrak{m}$.
 If it is $1$, we have $e(x_{1})=1$ mod.
-$\mathsf{m}R[x_{1}]$ and we can take $u_{1}=1$.
+$\mathfrak{m}R[x_{1}]$ and we can take $u_{1}=1$.
 We can thus assume $e(x_{1})\cdots e(x_{n})=0$ mod.
-$\mathsf{m}$.
+$\mathfrak{m}$.
 We write $e_{I}=\prod_{i\in I}e(x_{i})$ for $I$ subset of $1,\dots,n$.
 We let [^3] The $e_{I},~|I|=k$, are pairwise conjugate by the action of the symmetric group and equality is decidable mod.
-$\mathsf{m}$.
+$\mathfrak{m}L$.
 $k<n$ be such that $e_{I}\neq 0$ mod.
-$\mathsf{m}$ if $|I|=k$ and $e_{I}=0$ mod.
-$\mathsf{m}$ if $|I|=k+1$.
+$\mathfrak{m}L$ if $|I|=k$ and $e_{I}=0$ mod.
+$\mathfrak{m}L$ if $|I|=k+1$.
 Note that the $e_{I}$ for $|I|=k$ are pairwise orthogonal mod.
-$\mathsf{m}$.
+$\mathfrak{m}L$.
 We have $\Sigma_{|I|=k}e_{I}=1$ mod.
-$\mathsf{m}$ since it is an element in $R$, which is idempotent mod.
-$\mathsf{m}$.
-Hence the family of the $e_{I},~|I|=k$, forms an FSOI [^4] A “Fundamental System of Orthogonal Idempotents” ([LQ, Sec.
+$\mathfrak{m}$ since it is an element in $R$ which is idempotent mod.
+$\mathfrak{m}$.
+Hence the family of the $e_{I},~|I|=k$, forms an FSOI [^4] A “Fundamental System of Orthogonal Idempotents” ([LQ, Sect.
 II-4]).
 mod.
-$\mathsf{m}A$.
+$\mathfrak{m}L$.
 We deduce that
-$Q(X)=\prod_{|I|=k}(X-e_{I})\text{,}$
-which is a monic polynomial in $R[X]$, is of the form $X^{N-1}(X-1)$ mod. $\mathsf{m}R[X]$ with $N=\binom{n}{k}$.
-Since $R$ is Henselian, we have $\alpha$ in $R$ such that $Q(\alpha)=0$ and $\lambda=Q^{\prime}(\alpha)$ is invertible. If we write $Q_{I}=Q/(X-e_{I})$ then $Q^{\prime}(X)=\Sigma_{I}Q_{I}$. We take
-$v_{I}=\frac{1}{\lambda}Q_{I}(\alpha)$
+$g(X)=\prod_{|I|=k}(X-e_{I})\text{,}$
+which is a monic polynomial in $R[X]$, is of the form $X^{N-1}(X-1)$ mod. $\mathfrak{m}R[X]$ with $N=\binom{n}{k}$.
+Since $R$ is Henselian, we have $\alpha$ in $R$, $\alpha=1$ mod.
+$\mathfrak{m}$, such that $g(\alpha)=0$; furthermore $\lambda=g^{\prime}(\alpha)$ is invertible with $\lambda=1$ mod.
+$\mathfrak{m}$.
+If we write $g_{I}=g/(X-e_{I})$ then $g^{\prime}(X)=\Sigma_{I}g_{I}$.
+We take
+$v_{I}=\frac{1}{\lambda}g_{I}(\alpha)$
 and we then have that the $v_{I},~|I|=k$, form an FSOI in $R[x_{1},\dots,x_{n}]$.
 We have $v_{I}=\Pi_{J\neq I}(1-e_{J})=e_{I}$ mod.
-$\mathsf{m}A$.
-Let $L_{1}$ be the set of all $I$ such that $1$ is in $I$ and $|I|=k$.
-If we write $u_{1}$ for $\Sigma_{I\in L_{1}}v_{I}$, we have $u_{1}$ in $R[x_{1}]$ and $u_{1}$ is equal to $\Sigma_{I\in L_{1}}e_{I}=e(x_{1})$ mod.
-$\mathsf{m}A$.
-This equality follows from $e(x_{1})=e(x_{1})(\Sigma_{I}e_{I})=\Sigma_{I}e(x_{1})e_{I}=\Sigma_{I\in L_{1}}e_{I}$ mod.
-$\mathsf{m}A$.
-Hence $e(x_{1})-u_{1}\in R[x_{1}]\cap\mathsf{m}A=\mathsf{m}R[x_{1}]$.
+$\mathfrak{m}L$.
+Let $\mathcal{I}_{1}$ be the set of all $I$ such that $1$ is in $I$ and $|I|=k$.
+If we write $u_{1}$ for $\Sigma_{I\in\mathcal{I}_{1}}v_{I}$, we have $u_{1}$ in $R[x_{1}]$ and $u_{1}$ is equal to $\Sigma_{I\in\mathcal{I}_{1}}e_{I}=e(x_{1})$ mod.
+$\mathfrak{m}L$.
+This equality follows from $e(x_{1})=e(x_{1})(\Sigma_{I}e_{I})=\Sigma_{I}e(x_{1})e_{I}=\Sigma_{I\in\mathcal{I}_{1}}e_{I}$ mod.
+$\mathfrak{m}L$.
+Hence $e(x_{1})-u_{1}\in R[x_{1}]\cap\mathfrak{m}L=\mathfrak{m}R[x_{1}]$.
 ∎
-For completeness, we recall the statement and proof of Nakayama’s Lemma [LQ, Lemma IX-2.1].
-Lemma 2.2.
-Let $M$ be a finitely generated $R$-module and $N$ an $R$-submodule such that $M=N+\mathsf{m}M$, then $M=N$.
+Theorem 2.1 ([ALP, Theorem 5.9]).
+If $A$ is a commutative $R$-algebra which is finitely generated as an $R$-module and $a$ in $A$ is idempotent mod.
+$\mathfrak{m}A$, then we can find $e$ in $A$ idempotent in $A$ and such that $e=a$ mod.
+$\mathfrak{m}A$.
+The following generalisation is not in [ALP] and will be crucial in the proof of Lemma 3.1.
+Corollary 2.1.
+If $A$ is a (not necessarily commutative) $R$-algebra which is finitely generated as an $R$-module and $a$ in $A$ is idempotent mod.
+$\mathfrak{m}A$, then we can find $e$ in $A$ idempotent in $A$ and such that $e=a$ mod.
+$\mathfrak{m}A$.
 Proof.
-Let $u$ be a vector column of generators of $M$: then we can find a vector $v$ in $N$ and a matrix $P$ with elements in $\mathsf{m}$ such that $u=v+Pu$.
-We then have (IP)u=v(I-P)u=v and hence the elements of $v$ are generators of $M$ since $\det(I-P)$ is invertible.
-∎
-Corollary 2.1.
-Let $A$ be an $R$-algebra which is finitely generated as an $R$-module, and $u$ in $A$ which is invertible mod. $\mathsf{m}A$. Then $u$ is invertible in $A$.
-Proof.
-We have $A=uA+\mathsf{m}A$ and we can apply the previous Lemma. ∎
-Lemma 2.3 ([ALP, Prop. 5.8]).
-Let $P$ in $R[X]$ be a monic polynomial in $R[X]$ and assume that we have a decomposition $P=fg$ mod.
-$\mathsf{m}R[X]$ with $f$ and $g$ monic polynomials coprime mod.
-$\mathsf{m}$.
-Then we can find a decomposition $P=FG$ in $R[X]$ with $F=f$ and $G=g$ mod.
-$\mathsf{m}$.
-Proof.
-We let $R[x]=R[X]/(P)$.
-We can find an idempotent $e$ in $R[x]$ mod.
-$\mathsf{m}R[x]$ such that (f)=(e),(g)=(1e)(f)=(e),~(g)=(1-e) mod.
-$\mathsf{m}R[x]$ [ALP, Prop.
-3.11].
-Using Lemma 2.1, we may assume $e$ idempotent in $R[x]$.
-Let $n$ (resp.
-$m$) be the degree of $f$ (resp.
-$g$).
-Then $e,ex,\dots,ex^{m-1}$ and $(1-e),(1-e)x,\dots,(1-e)x^{n-1}$ generate $R[x]$ as an $R$-module mod.
-$\mathsf{m}$.
-By Lemma 2.2, they generate the $R$-module $R[x]$.
-In particular, $ex^{m}$ belongs to the $R$-module generated by $e,ex,\dots,ex^{m-1}$ (since $e$ is idempotent in $R[x]$) and we find in this way $G$ monic of degree $m$ such that $eG(x)=0$ in $R[x]$.
-Similarly we find $F$ monic of degree $n$ such that (1e)F(x)=0(1-e)F(x)=0 in $R[x]$.
-We then have $F(x)G(x)=F(x)eG(x)=0$ and hence $P$ divides $FG$.
-But they have the same degree, hence $P=FG$.
-∎
-Theorem 2.1 ([ALP, Thm. 5.9]).
-If $A$ is an $R$-algebra which is finitely generated as an $R$-module and $a$ in $A$ is idempotent mod.
-$\mathsf{m}A$, then we can find $e$ in $A$ idempotent in $A$ and such that $e=a$ mod.
-$\mathsf{m}A$.
-Proof.
-We consider the subalgebra $R[a]$ of $A$.
-Since $A$ is finitely generated as an $R$-module we first find $F$ monic in $R[X]$ such that $F(a)=0$.
-We decompose $F=X^{n}(1-X)^{m}h$ mod.
-$\mathsf{m}R[X]$, with $h$ monic and $h(0)$, $h(1)$ not in $\mathsf{m}$.
-If $n=0$ we have $F(0)=h(0)$ mod.
-$\mathsf{m}$.
-Since $F(a)=0$, we get that $a$ is invertible mod.
-$\mathsf{m}A$, and hence invertible in $A$ using Corollary 2.1, and we can take $e=1$.
-Similarly, if $m=0$, we get that 1a1-a is invertible and we can take $e=0$.
-In the remaining case, we can find using Lemma 2.3 a decomposition in monic polynomials $F=PQH$ in $R[X]$ with $P=X^{n},~Q=(1-X)^{m},~H=h$ mod.
-$\mathsf{m}$.
-We also have $U$ and $V$ in $R[X]$ such that X(1X)U+HV=Res(X(1X),H)=1X(1-X)U+HV=Res(X(1-X),H)=1 [LQ, Fact 3-7.1], [ALP, Lemma 3.10], and so $H(a)V(a)=1$ mod.
-$\mathsf{m}A$, and so $H(a)$ is invertible mod.
-$\mathsf{m}A$ and so invertible in $A$ using Corollary 2.1.
-Since $F(a)=P(a)Q(a)H(a)=0$, we have $P(a)Q(a)=0$ in $A$.
-We also have $P(a)=a$ and Q(a)=1aQ(a)=1-a mod.
-$\mathsf{m}A$ and so $P(a)+Q(a)=1$ mod.
-$\mathsf{m}A$ and is also invertible in $A$, by Corollary 2.1, with inverse $\mu$.
-We can then take $e=\mu P(a)$.
+Write $\varphi(x)=x^{2}-x$ and consider the finitely generated algebra $B=R[a]$.
+Since $\varphi(a)=a^{2}-a$ is in $\mathfrak{m}A$, the constant term of its characteristic polynomial is in $\mathfrak{m}A\cap R=\mathfrak{m}$, so that we have $N$ such that $\varphi(a)^{N}$ is in $\mathfrak{m}B$.
+We can then consider the following Newton quadratic iteration of $a$ inside $B$ by defining inductively $u_{0}=a$ and $u_{n+1}=u_{n}-(2u_{n}-1)\varphi(u_{n})$ (compare [LQ, Corollary III.10.4]).
+One checks that $u_{n}=a$ mod.
+$\mathfrak{m}A$ and that $\varphi(u_{n})$ is in $\varphi(u_{n-1})^{2}B$, so that it is in $\mathfrak{m}B$ for $n$ big enough, i.e.
+$u_{n}$ is idempotent mod.
+$\mathfrak{m}B$.
+We can now use Theorem 2.1 for the commutative $R$-algebra $B$ to lift $u_{n}$ to an idempotent $e$ in $B$ such that $e=u_{n}$ mod.
+$\mathfrak{m}B$.
+We also have $e=a$ mod.
+$\mathfrak{m}A$.
 ∎
@@
   1. 1.
-    AA is projective of finite type;
+    AA is finitely generated projective;
   2. 2.
-    the canonical map $A\otimes_{R}A^{op}\rightarrow End_{k}(A)$ is an isomorphism.
+    the canonical map $A\otimes_{k}A^{\mathrm{op}}\rightarrow\operatorname{End}_{k}(A)$ is an isomorphism.
 Note that we allow the trivial algebra to be an Azumaya algebra. Note also that any matrix algebra is Azumaya.
-One can equivalently describe the algebra structure by a bilinear map $\mu:A\times A\rightarrow A$; its scalar extension along $k\rightarrow S$ is the unique $S$-algebra $A_{S}$ through which every bilinear map extending $\mu$ factors.
+In this section $k$ will denote a commutative base ring, and $R$ will denote the “generic” separably closed commutative $k$-algebra, model of the coherent theory $T_{1}$ described in the first section.
+One can equivalently describe the algebra structure by a bilinear map $\mu\colon A\times A\rightarrow A$; its scalar extension along $k\rightarrow S$ is the unique $S$-algebra $A_{S}$ through which every bilinear map extending $\mu$ factors.
 Thus all subsequent arguments can be phrased without explicit tensor notation.
-Over a finite free $k$-module, to give an algebra structure is represented concretely by a multiplication table.
-A module $A$, projective of finite type over $k$, is given concretely by a projection matrix over $k$ [LQ], and we can similarly represent the algebra structure by a multiplication table (so $m^{3}$ elements of $k$ if the projection matrix has finite size $m$).
-We can then compute [LQ, X, Theorem 1.6] comaximal elements $s_{1},\dots,s_{l}$ such that $A_{s_{i}}$ is free over $k_{s_{i}}$, and the second condition for being an Azumaya algebra becomes then that some determinants $\delta_{1},\dots,\delta_{l}$ are invertible [^5] After localisation to $s_{j}$ the Azumaya algebra becomes free of rank $r_{i}=n_{i}^{2}\leqslant m$..
-From this discussion follows that if $A$ is an Azumaya algebra, given by a multiplication table, and we have a $k$-algebra $k\rightarrow S$ then the algebra $A_{S}$ over $S$, with the same multiplication in $S$, is also an Azumaya algebra.
-We have the following characterisation, which generalizes the result that an algebra is central simple over a field if, and only if, it can be split by a separable extension of this field.
+Over a finitely generated free $k$-module, an algebra structure is given concretely by a multiplication table.
+A module $A$, finitely generated projective over $k$, is given concretely by a projection matrix over $k$ [LQ, Chap.
+I], and we can similarly represent the algebra structure by a multiplication table (so $m^{3}$ elements of $k$ if the projection matrix has finite size $m$).
+We can then compute [LQ, Theorem X-1.6] comaximal elements $s_{1},\dots,s_{l}$ such that $A[1/s_{i}]$ is free over $k[1/s_{i}]$, and the second condition for being an Azumaya algebra becomes then that $l$ determinants are invertible.
+[^5] After localisation to $s_{i}$ the Azumaya algebra becomes free of rank $r_{i}=n_{i}^{2}\leqslant m$.
+From this discussion follows that if $A$ is an Azumaya algebra, given by a multiplication table, and we have a commutative $k$-algebra $k\rightarrow S$, then the algebra $A_{S}$ over $S$, with the same multiplication table, is also an Azumaya algebra.
+We have the following characterisation, which generalises the result that an algebra is central simple over a field if, and only if, it can be split by a separable extension of this field.
 Theorem 3.1.
-AA is an Azumaya algebra over $k$ if, and only if, we can build a finite tree with root $k$ using the following, where at all leaves all scalar extensions $A_{S}$ are matrix algebras:
+AA is an Azumaya algebra over $k$ if, and only if, we can build a finite covering tree rooted at $k$ using the following, where at every leave the scalar extension $A_{S}$ is a matrix algebra:
   1. 1.
-    $S$ is covered by $S\rightarrow S[1/u_{i}]$ if $1=(u_{1},\dots,u_{n})$;
+    $S$ is covered by $\rightarrow S[1/s_{1}], …, SS$$\rightarrow S[1/s_{n}$] if $1=(s_{1},\dots,s_{n})$;
   2. 2.
-    $S$ is covered by $S\rightarrow S[X]/(P)$ if $P$ is monic and unramifiable.
-We prove first that if we have such a tree then the algebra $A$ is Azumaya.
-If we have such a tree starting from $k$, notice that each extension $S\rightarrow S[X]/(P)$ for $P$ monic is faithfully flat, and $S\rightarrow\prod_{i}S[1/u_{i}]$ if $1=(u_{1},\dots,u_{n})$ is also faithfully flat.
-It follows then that $A$ is finite projective [LQ, Thm.
-VIII-6.7].
-The map $A\otimes_{k}A^{op}\rightarrow End_{k}(A)$ is then a bijection since it becomes a bijection by a faithfully flat extension.
+    $S$ is covered by $S\rightarrow S[X]/(f)$ if $f$ is monic and unramifiable.
+We shall use the local-global principle and the descent principle with respect to faithfully flat extension for projective modules ([LQ, Theorem VIII-6.7]).
+For completeness, we recall the statement.
+Note that the proof in [LQ] is constructive, using the concrete definition of faithfully flat extension and the fact that, also constructively, a module is finitely generated projective if, and only if, it is finitely presented and flat.
+Theorem 3.2 (Descent).
+Let $A\rightarrow B$ be a faithfully flat extension and $M$ an $A$-module.
+If $M\otimes_{A}B$ is finitely generated projective, then so is $M$.
+If we have $1=(s_{1},\dots,s_{n})$ in $A$ and each $M[1/s_{i}]$ is finitely presented and projective over $A[1/s_{i}]$ then so is $M$.
+We prove first that if we have such a covering tree then the algebra $A$ is Azumaya.
+If we have such a covering tree rooted at $k$, notice that each extension $S\rightarrow S[X]/(f)$ for $f$ monic is faithfully flat, and $S\rightarrow\prod_{i}S[1/s_{i}]$ is also faithfully flat if $1=(s_{1},\dots,s_{n})$.
+It follows then that $A$ is finitely generated projective by Theorem 3.2.
+The map $A\otimes_{k}A^{\mathrm{op}}\rightarrow\operatorname{End}_{k}(A)$ is then a bijection since it becomes a bijection by a faithfully flat extension.
 (As we saw before, to be a bijection is witnessed by the fact that some determinants are invertible, and an element is invertible if it becomes invertible by a faithfully flat extension.)
 For the converse, we look at the algebra $A$ (given concretely by its multiplication table) inside the étale topos over $k$, using the site defined in the first section.
-(This is similar to the constructive use of the algebraic closure in [MC].) The generic ring $R$, defined by $R(S)=S$ for $S$ a finitely presented $k$-algebra, is, in the internal logic of this topos, a local strictly Henselian ring.
-The algebra $A$ defines then an $R$-algebra $A_{R}(S)=A_{S}$, which is an Azumaya algebra, given by the same multiplication table as the one of $A$.
-If we can prove in intuitionistic logic that an Azumaya algebra over a strictly Henselian ring is a matrix algebra, we will get, using the interpretation of intuitionistic logic in a sheaf model [Joh, MC], a covering tree of $k$ which gives Theorem 3.1.
-Internally in the classifying topos of stricly Henselian $k$-algebras, the object $R$ is local strictly Henselian and the claim that $A_{R}$ is a matrix algebra is a coherent statement, provable intuitionistically.
-Such a proof yields externally a finite covering tree of $k$ satisfying the conditions above.
+(This is similar to the constructive use of the algebraic closure in [MC].) The generic ring $R$, defined by $R(S)=S$ for $S$ a finitely presented commutative $k$-algebra, is, in the internal logic of this topos, a separably closed local ring.
+The algebra $A$ defines then an $R$-algebra $A_{R}(S)=A_{S}$, which is the Azumaya algebra given by the same multiplication table as the one for $A$.
+If we can prove in intuitionistic logic that an Azumaya algebra over a separably closed ring is a matrix algebra, we will get, using the interpretation of intuitionistic logic in a sheaf model [Joh, MC], a covering tree rooted at $k$ which yields Theorem 3.1.
+Internally in the classifying topos of separably closed local commutative $k$-algebras, the object $R$ is separably closed local and the claim that $A_{R}$ is a matrix algebra is a coherent statement, provable in intuitionistic logic.
+Such a proof yields externally a finite covering tree rooted at $k$ satisfying the conditions above.
 The converse is a consequence of the following two results.
 Lemma 3.1.
-If $A$ is an Azumaya algebra over $R$, which is a local and residually discrete and separably closed $k$-algebra, then $A$ is a matrix algebra.
+If $A$ is an Azumaya algebra over $R$, which is a residually discrete separably closed local commutative $k$-algebra, then $A$ is a matrix algebra over $R$.
 Proof.
-We know that $R$ is Henselian and that $R/\mathsf{m}$ is separably closed by Theorem 1.1.
-We simplify the argument in Milne’s course notes ([Milne, Chapter IV, Proposition 1.6]), using the constructive development in [ALP] and [CLN].
-If $\mathsf{m}$ is the maximal ideal of $R$, we have that $A$ is split over $R/\mathsf{m}$ ([CLN]), and hence we have a matrix algebra decomposition $e_{ij}$ mod.
-$\mathsf{m}A$.
-We can then lift the idempotent $e_{11}$ mod.
-$\mathsf{m}A$ to some idempotent $e$ in $R$ using Theorem 2.1.
-We have that $Ae$ is projective and hence free since $R$ is a local ring.
-The map $A\rightarrow End_{R}(Ae)$ is then an isomorphism, since it is a map between two free $R$-modules of same dimension, and it is residually an isomorphism.
+We know that $R$ is Henselian and that $R/\mathfrak{m}$ is separably closed by Theorem 1.1.
+We simplify (and correct slightly) the argument in Milne’s course notes ([Milne, Chapter IV, Proposition 1.6]), using the constructive development in [ALP] and [CLN].
+If $\mathfrak{m}$ is the maximal ideal of $R$, we have that $A$ is split over $R/\mathfrak{m}$ ([CLN]), and hence we have a matrix algebra decomposition $e_{ij}$ mod.
+$\mathfrak{m}A$.
+We lift the idempotent $e_{11}$ mod.
+$\mathfrak{m}A$ to some idempotent $e$ in $A$ using Corollary 2.1.
+The $R$-module AeAe is finitely generated projective and hence free since $R$ is a local ring.
+The map $A\rightarrow\operatorname{End}_{R}(Ae)$ is then an isomorphism, since it is a map between two free $R$-modules of same dimension, and it is residually an isomorphism.
 ∎
-The next proof uses Barr’s Theorem, similar to the way it is used in [Penon, Wraith], but the difference is that we only need a constructively valid version of this theorem.
+The next proof uses Barr’s Theorem in a way similar to [Penon, Wraith], but the difference is that we only need a constructively valid version of this theorem.
 Corollary 3.1.
-If $A$ is an Azumaya algebra over $R$, which is a local and separably closed $k$-algebra, then $A$ is a matrix algebra.
+If $A$ is an Azumaya algebra over $R$, which is a separably closed local commutative $k$-algebra, then $A$ is a matrix algebra over $R$.
 Proof.
@@
 ∎
-Theorem 3.1 follows by interpreting this result in the classifying topos of the theory of strictly Henselian $k$-algebras.
+Theorem 3.1 follows by interpreting this result in the classifying topos of the theory of separably closed local commutative $k$-algebras.
 The statement of Theorem 3.1 can be seen as a generalisation of the fact that over a field, a central simple algebra becomes a matrix algebra by a faithfully flat extension [CLN] (classically over an algebraic extension).
-We can use such a tree to give arguments by “tree induction”. We give some simple examples.
+We can use such a covering tree to give arguments by “tree induction”. We give some simple examples.
 Lemma 3.2.
@@
 Proof.
-This means that if $a$ in $A$ satisfies $ax=xa$ for all $x$ in $A$ then $a$ is in $R$.
-We prove this by “tree induction” using Theorem 3.1.
+This means that if $a$ in $A$ satisfies $ax=xa$ for all $x$ in $A$ then $a$ is in $k$.
+We prove this by tree induction using Theorem 3.1.
 The statement holds on the leaves (it holds for a matrix algebra) and descends to the root.
@@
 Lemma 3.3.
-If $R$ is a local ring and $A$ a matrix algebra of rank $\geqslant 1$ [LQ, Def. X-2.2] then any automorphism of $A$ is of the form $x\mapsto axa^{-1}$ for some $a$ in $GL_{n}(A)$.
+If $R$ is a local ring and $A$ a matrix algebra of rank $\geqslant 1$ [LQ, Definition X-2.2] then any automorphism of $A$ is of the form $x\mapsto axa^{-1}$ for some $a$ in $\operatorname{GL}_{n}(A)$.
 Proof.
-Let $A=End_{R}(R^{n})$ with $v_{1},\dots,v_{n}$ the canonical basis of $R^{n}$ and define $e_{ij}v_{k}=\delta_{jk}v_{i}$.
+Let $A=\operatorname{End}_{R}(R^{n})$ with $v_{1},\dots,v_{n}$ the canonical basis of $R^{n}$ and define $e_{ij}v_{k}=\delta_{jk}v_{i}$.
 Let $\psi$ be an automorphism of $A$.
@@
 Let $w_{1}$ be a basis of $V_{1}$ and $w_{j}=p_{j1}w_{1}$.
-If $a(v_{i})=w_{i}$ we have $\psi(x)=axa^{-1}$.
+If $a(v_{i})=w_{i}$ then we have $\psi(x)=axa^{-1}$.
 ∎
-For the next result, we use the local-global principle and the descent principle with respect to faithfully flat extension for projective modules ([LQ, Theorem VIII-6.7]).
-For completeness, we recall the statement.
-Note that the proof in [LQ] is constructive, using the concrete definition of faithfully flat extension and the fact that, also constructively, a module is finite projective if, and only if, it is finitely presented and flat.
-Theorem 3.2 (Descent).
-Let $A\rightarrow B$ be a faithfully flat extension and $M$ an $A$-module.
-If $M\otimes_{A}B$ is finite projective, then so is $M$.
-If we have $1=(s_{1},\dots,s_{n})$ in $A$ and each $M_{s_{i}}$ is finite projective over $A_{s_{i}}$ then so is $M$.
 Theorem 3.3 (Skolem-Noether).
@@
 If $A$ is an Azumaya algebra, of rank $\geqslant 1$ as a projective module over a ring $k$, and $\psi$ is an automorphism of $A$, then
-$M=\{a\in A~|~\forall_{x}\,ax=\psi(x)a\}$ and $N=\{a\in A~|~\forall_{x}\,a\psi(x)=xa\}$
+$M=\{a\in A~|~\forall_{x}\,ax=\psi(x)a\}\quad\text{and}\quad N=\{a\in A~|~\forall_{x}\,a\psi(x)=xa\}$
 are projective $k$-modules of rank $1$ and $M\otimes_{k}N\simeq k$ and $MN=k1$. Furthermore, $M$ and $N$ are direct factors in $A$.
@@
 We have on the leaves $M_{S}N_{S}\subseteq S1$ and this statement descends to the root by tree induction.
-This can be used to define a map $M\rightarrow Hom(N,k),~x\mapsto(y\mapsto xy)$.
-This map is an isomorphism on the leaves, and this also descends to the root by tree induction.
+This can be used to define a map $M\rightarrow\operatorname{Hom}(N,k),~x\mapsto(y\mapsto xy)$.
+This map is an isomorphism on the leaves, and this statement also descends to the root by tree induction.
 Finally, $M$ and $N$ are direct factors on the leaves, and this fact descends as well.
@@
 To get the theory of algebraically closed local rings, we add to the theory of local rings the axioms
-$\exists_{x}~P(x)=0$
-for any monic nonconstant polynomial $P$.
-Wraith conjectured in [Wraith] that this should define the classifying theory of the fppf topos. This question is discussed further in Blechschmidt’s PhD thesis [BT].
+$\exists_{x}~f(x)=0$
+for any monic nonconstant polynomial $f$.
+Wraith conjectured in [Wraith] that this should define the classifying theory of the fppf topos. This question is discussed further in Blechschmidt’s Ph.D. thesis [BT].
 The same argument as in Lemma 3.1 shows the following result.
 Lemma 4.1.
-If $A$ is an Azumaya algebra over $R$, which is a local and residually discrete and algebraically closed $k$-algebra, then $A$ is a matrix algebra.
+If $A$ is an Azumaya algebra over $R$, which is a residually discrete algebraically closed local commutative $k$-algebra, then $A$ is a matrix algebra.
 We can then prove as in the previous section
 Theorem 4.1.
-AA is an Azumaya algebra over $k$ if, and only if, we can build a finite tree of root $k$ using the following, where all leaves are matrix algebras:
+AA is an Azumaya algebra over $k$ if, and only if, we can build a finite covering tree rooted at $k$ using the following, where all leaves are matrix algebras:
   1. 1.
-    $S$ is covered by $S\rightarrow S[1/u_{i}]$ if $1=(u_{1},\dots,u_{n})$;
+    $S$ is covered by $\rightarrow S[1/s_{1}], …, SS$$\rightarrow S[1/s_{n}$] if $1=(s_{1},\dots,s_{n})$;
   2. 2.
-    $S$ is covered by $S\rightarrow S[X]/(P)$ if $P$ is monic nonconstant.
+    $S$ is covered by $S\rightarrow S[X]/(f)$ if $f$ is monic nonconstant.
 5 Conclusion
 This paper is a beginning of the theory of Azumaya algebras in a constructive setting.
-The natural next step will be to analyse Gabber’s Theorem [KO] that the group of Azumaya algebras is the torsion part of the second cohomology group with coefficients in $GL_{n}(R)$ for the étale topology.
+The natural next step will be to analyse Gabber’s Theorem [KO] that the group of Azumaya algebras is the torsion part of the second cohomology group with coefficients in $\operatorname{GL}_{n}(k)$ for the étale topology.